### v1

# Adapt Tool Schemas to the Models

Abstract
Small language models (SLMs) offer significant computational advantages for tool-augmented AI systems, yet they struggle with tool-use tasks, particularly in selecting appropriate tools and identifying correct parameters. A common failure mode is schema misalignment: models hallucinate plausible but non-existent tool names that reflect naming conventions internalized during pretraining but absent from the provided tool schema. Rather than forcing models to adapt to arbitrary schemas, we propose adapting schemas to align with models’ pretrained knowledge. We introduce PA-Tool (Pretraining-Aligned Tool Schema Generation), a training-free method that leverages peakedness—a signal from contamination detection indicating pretraining familiarity—to automatically rename tool components. By generating multiple candidates and selecting those with highest output concentration across samples, PA-Tool identifies pretrain-aligned naming patterns. Experiments on MetaTool and RoTBench show improvements of up to 17% points, with schema misalignment errors reduced by 80%. PA-Tool enables small models to approach state-of-the-art performance while maintaining computational efficiency for adaptation to new tools without retraining. Our work demonstrates that schema-level interventions can unlock the tool-use potential of resource-efficient models by adapting schemas to models rather than models to schemas.  [^1] We will release our code upon publication of the paper.
Don’t Adapt Small Language Models for Tools;
Adapt Tool Schemas to the Models
Jonggeun Lee, Woojung Song, Jongwook Han, Haesung Pyun, Yohan Jo Graduate School of Data Science, Seoul National University {jonggeun.lee, opusdeisong, johnhan00, haesung.pyun, yohan.jo}@snu.ac.kr
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
Tool-augmented language models have become essential components of modern AI systems, with tool-agent frameworks being widely adopted across various applications Schick et al. (2023). As these systems mature, there is growing interest in deploying small language models (SLMs) due to computational constraints and the need for edge deployment Belcak et al. (2025); Chen et al. (2025b). In these systems, SLMs must perform two critical operations: tool selection (identifying which function to call) and parameter identification (providing appropriate arguments). However, small models struggle significantly with these tasks, exhibiting severe performance degradation particularly for models below 10B parameters that are most viable for edge deployment Patil et al. (2025); Erdogan et al. (2024).
Figure 1: Effect of schema alignment on tool invocation. Top: Models learn tool documentation during pretraining. Bottom-Left: When schemas misalign with these internalized patterns, models generate plausible but non-existent tools. Bottom-Right: Schema alignment with pretrained knowledge prevents such errors.
A common failure pattern is schema misalignment: models hallucinate plausible but nonexistent components within the tool schema—a hierarchical structure consisting of tools and parameters, each with associated descriptions. As illustrated in Figure 1 (Top), models encounter various tool documentation patterns during pretraining, internalizing common naming conventions and API structures. However, when deployed with a specific tool schema, misalignment can occur. For instance (Figure 1, Bottom-Left), when the actual tool name to invoke is get_user_id, the model may instead generate get_customer_id—a similar but non-existent tool within the tool schema—leading to execution failure. Critically, these misaligned names are not random; they reflect naming patterns the model frequently encountered during pretraining but that are absent from the current tool schema. This raises a fundamental question: rather than forcing models to adapt to arbitrary tool schemas, what if we adapt the schemas to align with the naming conventions the models have internalized deeply? (Figure 1, Bottom-Right)
We hypothesize that by renaming tools to match patterns in the model’s pretraining data—transforming get_user_id to get_customer_id if the latter better aligns with learned representations—we can reduce errors and improve tool-calling accuracy. To operationalize this idea, we propose PA-Tool (Pretraining-Aligned Tool Schema Generation), which generates a mapping between original tool names and pretrain-aligned alternatives. To identify such pretrain-aligned patterns, we leverage insights from contamination detection methods Dong et al. (2024), which detect whether models encountered specific data during pretraining. A key finding from these studies is that contaminated data—patterns frequently seen during training—exhibits peakedness: models generate highly similar outputs across multiple sampling attempts, creating a concentrated distribution. This occurs because repeated exposure to specific patterns during pretraining strengthens internal representations, reducing output variance around familiar patterns Shi et al. (2024a); Li (2023). We adopt this peakedness as a signal for the pretraining familiarity of tool schemas.
PA-Tool renames each component in a schema in three stages: (1) instructing the target language model to generate multiple candidate names based on a description of the component Chen et al. (2021), (2) computing the peakedness of each candidate, i.e., the number of other candidates that are sufficiently similar based on character-level edit distance Levenshtein (1965), and (3) selecting the candidate with the highest peakedness as the new name, as it is considered to best align with the model’s pretrained representation.
Experiments on MetaTool and RoTBench with SLMs demonstrate PA-Tool’s effectiveness across diverse scenarios. On MetaTool, PA-Tool achieves substantial improvements across all subtasks, with the most significant gains when models must determine when no suitable tool exists (up to 17.0%) and when reasoning about multiple tools simultaneously (up to 9.6%). These improvements demonstrate PA-Tool’s effectiveness in addressing schema misalignment, which becomes particularly critical in complex tool selection scenarios. On RoTBench, PA-Tool maintains robust performance in both single-turn (5-10%) and multi-turn settings (4-10%), demonstrating that alignment benefits persist across extended conversational contexts. Notably, PA-Tool enables SLMs to approach the performance of closed-source models, with configurations like Qwen2.5-7B-Instruct matching or exceeding Claude Sonnet 4.5 on tool selection tasks.
Our comprehensive analysis reveals that PA-Tool’s effectiveness stems from directly addressing the predominant failure mode in SLMs. Schema misalignment errors—where models generate plausible but non-existent tool names—decrease by 80.0% with PA-Tool, while other error types show more modest reductions (18.8-24.0%). This targeted impact demonstrates that aligning schemas with pretrained knowledge resolves a bottleneck that disproportionately affects small models. Furthermore, PA-Tool’s training-free nature makes it highly practical for real-world deployment. This method requires only a simple one-time schema mapping without any model modification, enabling a model to use diverse tool sets without costly and complex retraining, and without having to worry about catastrophic forgetting or overfitting. This schema-level intervention represents an alternative approach: rather than forcing models to conform to arbitrary documentation standards, we adapt the interface to match the models’ internalized knowledge, enabling efficient tool use while preserving the computational advantages of small models.
Our contributions are summarized as follows:
  • We propose PA-Tool, a training-free schema optimization method that repurposes peakedness—a contamination detection signal—to identify pretrain-aligned tool names.
  • We demonstrate improvements of up to 17% points across SLMs on MetaTool and RoTBench, with benefits extending from tool selection to parameter identification in both single-turn and multi-turn settings.
  • We demonstrate that through a simple, easily deployable mapping interface, small models can approach closed-source model performance in specific aspects while maintaining computational efficiency.
2 Related Works
2.1 Emerging Agentic Frameworks
Recent advances in LLMs have enabled sophisticated agentic frameworks that decompose complex tasks into specialized modules, each coordinated by dedicated agents Shinn et al. (2023); Madaan et al. (2023); Wu et al. (2025); Agashe et al. (2025); Sapkota et al. (2026). In these multi-agent systems, SLMs are increasingly replacing larger models within individual modules to reduce computational costs and latency while maintaining specialized functionalities Belcak et al. (2025); Cheng et al. (2024). However, these systems remain vulnerable to cascading failures when SLM agents malfunction in foundational tool interaction tasks such as tool selection and parameter identification Erdogan et al. (2024); Patil et al. (2025). A particularly challenging failure mode is schema misalignment: SLM errors often manifest as plausible yet incorrect outputs—such as generating tool names that seem reasonable but don’t exist in the actual schema—making them difficult to detect and correct through traditional validation mechanisms. Our work addresses this vulnerability by generating pretraining-aligned schemas that improve SLM reliability in these critical operations.
2.2 Improving Tool Utilization in LLMs
As LLMs have demonstrated the capability to interact with external tools Schick et al. (2023); Hsieh et al. (2023), research has focused on evaluating and improving their tool-use capabilities. Evaluation efforts have developed benchmarks ranging from fine-grained assessments of tool selection and parameter identification Huang et al. (2024); Li et al. (2023); Ye et al. (2024); Chen et al. (2024); Patil et al. (2025) to end-to-end multi-step evaluation Seo et al. (2025); Yao et al. (2025); Trivedi et al. (2024); Qin et al. (2024); Shim et al. (2025).
Approaches to improving tool-use capabilities follow two main directions. Training-based methods employ supervised fine-tuning Qin et al. (2024); Liu et al. (2025); Zhang et al. (2025) or reinforcement learning Qian et al. (2025); Shi et al. (2024b); Feng et al. (2025); Chen et al. (2025a), but they require substantial data or computational resources. Training-free methods refine tool documentation Yuan et al. (2025); Qu et al. (2025) or leverage interaction histories Wang et al. (2024); Cui et al. (2025); Zhao et al. (2024); Fu et al. (2024). However, existing training-free approaches primarily focus on improving descriptions rather than aligning the schema itself with model-preferred representations. Our work directly addresses this gap by generating schemas aligned with pretraining distributions, targeting the root cause of schema misalignments.
2.3 Data Contamination Detection
Data contamination, defined as overlap between pretraining data and evaluation benchmarks, can inflate performance through memorization. Early detection methods used n-gram overlap Brown et al. (2020), but this approach fails to detect semantic paraphrasing. Probability-based approaches such as Min-k% Prob Shi et al. (2024a) and perplexity analysis Li (2023) require token probabilities that are unavailable in closed-source models. LLM Decontaminator Yang et al. (2023) uses auxiliary models for semantic similarity but introduces additional dependencies. Most relevant to our work, CDD Dong et al. (2024) identifies memorized patterns by measuring peakedness in sampled candidates, making it applicable to any black-box model. We adapt this peakedness mechanism to identify pretrain-aligned schema patterns, transforming contamination detection into a constructive tool for schema optimization.
Figure 2: Overview of our PA-Tool framework.
3 PA-Tool : Pretraining-Aligned Tool Schema Generation
We now formalize our approach to generating pretrain-aligned tool schemas. A tool schema is hierarchical, structured documentation that defines tools and parameters along with detailed descriptions for each component. Our objective is to transform these descriptions into component names that the model is familiar with and that align with patterns the model has frequently encountered during pretraining. Specifically, we construct a dictionary mapping that associates each original component with a pretrain-aligned name—one that the language model would naturally generate based on its pretraining experience.
The core principle is straightforward: by analyzing the peakedness of a model’s output distribution when generating names from component descriptions, we can identify which naming patterns the model has most frequently encountered during pretraining. These pretrain-aligned names are hypothesized to improve the model’s ability to accurately invoke tools with correct parameters, reducing schema-related errors such as incorrect tool selections and parameter mismatches.
3.1 Framework Overview
Let $\mathcal{M}$ denote a pre-trained language model. Given a natural language description $d$ of a schema component’s functionality (e.g., a tool or a parameter description), our objective is to generate an optimal name $s^{*}$ that represents the naming pattern most deeply internalized by $\mathcal{M}$. As illustrated in Figure 2, our approach operates in three stages: (1) we provide $\mathcal{M}$ with a description of a component and instruct it to generate its name multiple times (Stage 1), (2) we compute peakedness scores that measure how many similar candidates cluster around each candidate name (Stage 2), and (3) we select the candidate with the highest peakedness as the pretrain-aligned name (Stage 3). By applying this process to each component in the schema hierarchy, we construct a dictionary mapping from original names to pretrain-aligned names. The complete algorithm is provided in Appendix A.
3.2 Stage 1: Candidate Generation
In this stage, we collect diverse candidate names of a component that the model may have encountered during the pretraining. As illustrated in Figure 2-1, given a component’s description, we sample $N$ candidate names $\mathcal{C}=\{s_{1},s_{2},\ldots,s_{N}\}$ from the language model with temperature $t\in(0,1]$. This temperature-controlled sampling explores the model’s learned distribution beyond the single greedy path, revealing diverse naming patterns Chen et al. (2021). The detailed prompt structure is provided in the Appendix B. Additionally, we generate a reference name $s_{\text{ref}}$ using greedy decoding ($t=0$) for tie-breaking purposes in the selection stage.
3.3 Stage 2: Peakedness Computation
Following contamination detection principles Dong et al. (2024), we analyze the local concentration of the output distribution to identify strongly memorized patterns. As illustrated in Figure 2-2, for each candidate name $s_{i}$, we compute its peakedness score by counting how many similar candidate names the model generates around it. The intuition is that candidate names with many similar variants indicate pretrain-frequent patterns that the model generates consistently, while isolated candidates suggest less familiar patterns.
To quantify this clustering behavior, we first define a similarity threshold $\tau$ based on the maximum character length in the candidate set:
$\tau=\alpha\cdot\ell_{\max}$ (1)
where $\ell_{\max}=\max_{s_{i}\in\mathcal{C}}|s_{i}|$ is the maximum character length across all candidates, and $\alpha\in[0,1]$ is a hyperparameter that controls the strictness of similarity. This length-adaptive threshold ensures that longer names are allowed proportionally more variation while maintaining consistent similarity criteria across different name lengths.
The peakedness score for each candidate $s_{i}$ is then computed as:
$\phi(s_{i})=\sum_{j\neq i}\mathbb{I}(d_{\text{edit}}(s_{i},s_{j})\leq\tau)$ (2)
where $d_{\text{edit}}(\cdot,\cdot)$ denotes the character-level edit distance using the Levenshtein algorithm Levenshtein (1965), and $\mathbb{I}(\cdot)$ is the indicator function that equals 1 when the condition is satisfied and 0 otherwise. This score counts the number of candidates that fall within the similarity threshold from $s_{i}$, reflecting how many similar names the model generates around this pattern.
3.4 Stage 3: Representative Name Selection
The representative name is selected as the candidate with the maximum peakedness:
$s^{*}=\underset{s_{i}\in\mathcal{C}}{\mathrm{arg\,max}}\;\phi(s_{i})$ (3)
This criterion identifies the naming pattern that the model generates most consistently across multiple sampling attempts, indicating it represents the most deeply internalized pattern from the training data. In case of ties where multiple candidates achieve the same maximum peakedness score, we break ties by selecting the candidate with the minimum edit distance to the reference name:
$s^{*}=\underset{s_{i}\in\mathcal{C}^{*}}{\mathrm{arg\,min}}\;d_{\text{edit}}(s_{i},s_{\text{ref}})$ (4)
where $\mathcal{C}^{*}=\{s_{i}\in\mathcal{C}:\phi(s_{i})=\max_{s_{j}\in\mathcal{C}}\phi(s_{j})\}$ contains all candidates with maximum peakedness.
This approach rests on the hypothesis that frequently occurring patterns in training data create local maxima in the model’s output distribution. When a naming pattern appears multiple times during training, the model develops stronger internal representations, leading to higher probability mass and reduced variance around these patterns. By identifying regions of high peakedness, we effectively locate these memorized naming conventions, which represent the most natural and well-formed component names according to the model’s learned knowledge. Through iterative application across all schema components (Figure 2-3), we obtain the final pretrain-aligned schema. An example of a schema generated through this process is provided in Appendix C.
4 Experimental Setup
Model Method MetaTool RoTBench
Tool Selection Single-turn Multi-turn
Similar Scenario Reliability Multi-tool Tool Sel. Param Iden. Tool Sel. Param Iden.
Closed-Source Models
GPT-4.1-mini Base 79.6 84.3 76.3 72.2 79.1 58.1 28.6 12.9
Gemini-2.5-Flash Base 70.0 79.8 89.2 77.3 82.9 56.2 22.9 12.9
Claude-Sonnet-4.5 Base 75.6 83.0 84.3 85.1 83.3 67.6 25.7 17.1
Open-Source Models
Qwen2.5-3B Base 48.7 55.3 83.6 75.1 1.9 0.0 1.4 0.0
Greedy 49.1 57.0 82.9 63.4 6.7 3.8 5.7 1.4
PA-Tool 50.0 58.8 86.2 72.6 6.7 1.9 7.1 4.3
Qwen2.5-7B Base 59.6 74.4 78.3 78.3 49.5 20.0 8.6 5.7
Greedy 60.6 75.4 85.4 82.7 50.5 21.0 14.3 8.6
PA-Tool 64.1 78.4 88.2 84.9 55.2 21.9 18.6 11.4
Llama3.2-3B Base 55.0 58.6 43.6 79.1 56.2 20.0 27.1 5.7
Greedy 57.7 58.9 39.8 70.4 59.1 20.0 28.6 4.3
PA-Tool 65.7 67.7 60.6 80.5 62.9 21.9 31.4 5.7
Llama3.1-8B Base 61.5 73.9 53.5 78.7 58.1 17.1 17.1 4.3
Greedy 64.6 72.9 51.5 78.9 63.8 18.1 17.1 4.3
PA-Tool 70.4 79.9 66.0 88.3 68.6 18.1 22.9 4.3
Table 1: Performance comparison on MetaTool and RoTBench. All metrics are reported as accuracy (%). Tool Sel. and Param Iden. denote Tool Selection and Parameter Identification, respectively. Bold indicates the best performance within each closed-source model and among all open-source models.
To evaluate whether PA-Tool improves models’ core tool-use capabilities, we conduct comprehensive experiments assessing two fundamental aspects: (1) Tool selection, whether models correctly identify the appropriate tool for a given query from available options, and (2) Parameter identification, whether models accurately extract necessary arguments to construct valid tool calls.
4.1 Benchmarks
MetaTool.
MetaTool (Huang et al., 2024) evaluates tool selection capabilities across 4,287 test cases in four subtasks: (1) Similar tests semantic comprehension by distinguishing tools with overlapping functionalities (e.g., Sudoku vs. Tic-Tac-Toe); (2) Scenario select appropriate tools based on user-specific contexts and requirements (e.g., software engineers, students); (3) Reliability assesses whether models can directly indicate when no suitable tool is available rather than hallucinating; and (4) Multi-tool measures whether models can correctly select multiple tools when tasks require composition of functionalities.
RoTBench.
RoTBench (Ye et al., 2024) evaluates whether models can robustly use tools when descriptions contain various noise perturbations (e.g. reversed names). However, we focus exclusively on the Clean environment to isolate the effects of PA-Tool on core tool-use capabilities rather than noise robustness. We evaluate on 175 test cases across 568 tools in both single-turn and multi-turn settings. RoTBench measures two fundamental tool use capabilities: (1) Tool Selection evaluates whether models correctly identify the appropriate tool; and (2) Parameter Identification assesses whether models accurately extract the required parameter set, conditioned on correct tool selection.
4.2 Models and Baselines
We focus on models with 8B parameters or fewer (SLMs), evaluating four open-source models: Qwen2.5-3B/7B (Qwen et al., 2025), Llama3.1-8B (Meta AI, 2024a), and Llama3.2-3B (Meta AI, 2024b). For each model, we evaluate three configurations: Base without any modifications, Greedy using greedy decoding for schema generation, and PA-Tool with our proposed method. To contextualize SLM performance and quantify PA-Tool’s improvements, we also compare against three state-of-the-art closed-source models: GPT-4.1-mini (OpenAI, 2024), Gemini 2.5 Flash (Basu Mallick et al., 2025), and Claude Sonnet 4.5 (Anthropic, 2025).
4.3 Implementation Details
When generating tool schemas with PA-Tool, we generate 32 candidates with temperature 0.4 and set $\alpha$ to 0.2. For all benchmark evaluations, we employ a one-shot prompting strategy where the prompt includes: (1) the task instruction, (2) the target tool description (not the complete schema), and (3) one demonstration example. We use greedy decoding to ensure reproducible results across all experiments. We use accuracy as the primary evaluation metric across both benchmarks, measuring the percentage of test cases where the model’s predictions exactly match the ground-truth labels. Detailed prompt templates are provided in Appendix B.
5 Main Results
We present our main experimental results demonstrating that PA-Tool consistently achieves the best performance across all models and tasks on both MetaTool and RoTBench benchmarks.
5.1 MetaTool: Consistent Improvements Across All Subtasks
Table 1 shows that PA-Tool consistently improves tool selection performance across all four MetaTool subtasks. In the Similar and Scenario tasks, PA-Tool achieves improvements of up to 10.7 % points, demonstrating that even simple alignment of tool names with the model’s internal knowledge effectively enhances tool selection. This phenomenon is more pronounced in the Multi-tool task, where gains reach up to 9.6 % points, with Llama3.1-8B improving from 78.7% to 88.3%. When tasks require identifying multiple tools simultaneously, the impact of schema misalignment compounds across each tool selection, making alignment particularly critical.
The Reliability task shows substantial improvements of up to 17.0%. For example, Llama3.2-3B improves from 43.6% to 60.6%. This task requires models to recognize when no suitable tool exists in the candidate list. Success depends on accurately understanding what each available tool does. PA-Tool aligns tool names with pretrained knowledge, enabling models to more clearly distinguish between available options and confidently determine when none match the query requirements.
Notably, the Greedy Schema baseline occasionally underperforms the Base model, with Llama3.2-3B showing 39.8% versus 43.6% in Reliability. This can be attributed to greedy decoding generating only a single schema candidate, which limits exploration of the schema space and increases susceptibility to suboptimal or misaligned schemas.
5.2 RoTBench: Persistent Benefits in Multi-Turn Interactions
RoTBench evaluates models in both single-turn and multi-turn conversational settings across two sequential stages: tool selection and parameter identification. In tool selection, PA-Tool demonstrates consistent improvements of 5-10% points across all models in single-turn scenarios. Importantly, these benefits remain robust in multi-turn settings with gains of 4-10% points in multi-turn settings, demonstrating that schema alignment remains effective across extended conversational contexts.
Parameter identification also shows improvements, though more modest than tool selection. The most notable gain appears in multi-turn settings, where Qwen2.5-7B nearly doubles from 5.7% to 11.4%. The smaller overall improvement stems from the nature of parameter naming. Parameters typically have explicit, descriptive names that already align well with common conventions (e.g., user_id with description "a unique user identifier"). Unlike tool names, which vary significantly across different tools, parameter names follow more standardized patterns. This leaves less room for alignment-based improvements.
5.3 Comparison with Closed-source Models
While our open-source models with PA-Tool show substantial improvements, a performance gap remains compared to state-of-the-art closed-source models. Claude Sonnet 4.5 achieves 83.3% tool selection and 67.6% parameter identification on RoTBench single-turn, significantly outperforming even our best configurations. However, PA-Tool narrows this gap, elevating SLMs closer to closed-source baselines while maintaining the computational efficiency advantages of SLMs.
Notably, in certain aspects of tool selection, PA-Tool enables small models to achieve competitive or superior performance. In MetaTool’s Multi-tool task, Llama3.1-8B with PA-Tool achieves 88.3%, surpassing all closed-source models including Claude Sonnet 4.5. Similarly, in the Reliability task, Qwen2.5-7B with PA-Tool reaches 88.2%, closely approaching Gemini 2.5 Flash’s 89.2%. These results demonstrate that schema alignment can enable resource-efficient models to match or exceed state-of-the-art systems in specific aspects of tool selection, particularly in tasks where schema misalignment is a primary failure mode.
6 Analysis
6.1 Error Analysis
Figure 3: Error type distribution for Llama3.1-8B on MetaTool tool selection tasks.
To understand how PA-Tool addresses different failure modes, we analyze error type distributions in Llama3.1-8B before and after applying PA-Tool. We categorize all incorrect predictions from both Base and PA-Tool configurations into three error types: Schema Misalignment Error, where the model generates non-existent or incorrectly named tools; Functional Confusion Error, where the model selects wrong tools with similar functionality; and Context Understanding Error, where the model selects functionally unrelated tools. Detailed error type definitions and experimental settings are provided in Appendix E.
Results. Figure 3 shows how PA-Tool affects different error types. Schema Misalignment Errors decrease dramatically by 80.0% (from 80 to 16 cases), confirming that aligning tool names with pretrained knowledge directly addresses this primary failure mode. In contrast, Functional Confusion Errors decrease by 24.0% (from 271 to 206 cases) and Context Understanding Errors decrease by 18.8% (from 607 to 493 cases). This stark difference—80.0% reduction for schema-related errors versus 18.8-24.0% for other types—demonstrates PA-Tool’s targeted impact. Schema alignment effectively eliminates naming-related failures, but distinguishing functionally similar tools and understanding complex query contexts require capabilities beyond schema optimization.
6.2 Integration with Supervised Fine-tuning
While PA-Tool is training-free, we demonstrate it remains effective when combined with supervised fine-tuning (SFT), addressing a key question: do models retain pretraining-aligned naming preferences after task-specific adaptation? Detailed information about SFT is presented in Appendix D. We compare four configurations: (1) Vanilla: base model with original schemas; (2) SFT only: fine-tuned model with original schemas; (3) PA-Tool only: base model with aligned schemas; (4) SFT + PA-Tool: SFT model with schemas aligned using the SFT model’s own PA-Tool generation.
Results. Table 2 shows PA-Tool provides consistent gains whether applied alone or with SFT. Notably, PA-Tool alone (third row) outperforms SFT alone (second row) across Similar, Reliability and Multi-tool by up to 10.1%, demonstrating superior efficiency without training. When combined, SFT + PA-Tool (fourth row) achieves the best performance on Similar (72.7%), Scenario (80.8%) and Multi-tool (89.9%) with consistent improvements over SFT alone across all subtasks.
Implications. The consistent gains demonstrate that pretraining-aligned naming preferences persist through fine-tuning. Schemas generated by PA-Tool using either the base or fine-tuned model are highly similar and both improve the model’s performance, indicating that task-specific adaptation does not override deeply internalized naming conventions from pretraining. This shows PA-Tool can deliver further performance gains on top of fine-tuning: schema-level alignment remains necessary to bridge the gap between arbitrary documentation standards and the model’s pretrained knowledge.
Configuration Similar Scenario Reliability Multi-tool
Llama3.1-8B 65.2 71.1 55.1 79.8
+ SFT 71.2 77.6 57.1 82.8
+ PA-Tool 72.2 75.9 67.2 89.9
+ SFT + PA-Tool 72.7 80.8 57.1 89.9
Table 2: Results of fine-tuning method and PA-Tool on MetaTool 10% random sampled test set.
6.3 Hyperparameter Analysis
We investigate the impact of PA-Tool’s three key hyperparameters on performance: the number of candidates ($N$), the similarity threshold ($\alpha$), and the sampling temperature ($t$). All experiments are conducted on the MetaTool benchmark, and we report the average accuracy across all four subtasks. Detailed results are provided in Appendix F.
Figure 4: Impact of hyperparameters on PA-Tool across different models. All results are averaged across four MetaTool subtasks. Top: Effect of the number of candidates (NN). Middle: Effect of similarity threshold ($\alpha$). Bottom: Effect of sampling temperature ($t$).
Number of candidates ($N$) (Figure 4, Top). Smaller models (3B) achieve stable performance with 16-32 candidates, while larger models (7-8B) require 32-64 candidates before performance plateaus. Beyond these ranges, additional candidates provide minimal gains.
Similarity threshold ($\alpha$) (Figure 4, Middle). Performance peaks at $\alpha=0.2$ consistently across all models. At $\alpha=0.1$, performance drops by 2-3 % points, while $\alpha=0.3$ shows similar degradation.
Sampling temperature ($t$) (Figure 4, Bottom). Performance remains stable across temperatures $t\in[0.2,1.0]$, varying within 1-2 % points. Moderate temperatures (t=0.4t=0.4-$0.6$) show slightly better results, though the differences are marginal.
7 Conclusion
We introduced PA-Tool, a training-free method that aligns tool schemas with models’ pretrained knowledge by leveraging peakedness as a pretraining familiarity signal. Experiments demonstrate improvements of up to 17% points with schema misalignment errors reduced by 80%, validating schema adaptation as an effective strategy for enhancing tool use in small language models. PA-Tool’s practical advantages make it particularly valuable for resource-constrained deployments—as a simple schema-level intervention, it can be applied without model training or fine-tuning, requiring only straightforward name mapping. By bridging pretrained knowledge and real-world tool interfaces, PA-Tool unlocks small models’ potential for tool-augmented applications while preserving computational efficiency.
Limitations
While PA-Tool effectively reduces schema misalignment errors by 80%, it shows more modest improvements for functional confusion and context understanding errors, indicating that schema optimization alone cannot address all failure modes. Our reliance on peakedness as an alignment signal assumes this metric reliably indicates pretraining familiarity. While validated across our experiments, this relationship may vary for models with substantially different training distributions. Additionally, our evaluation focuses on English-language schemas; the effectiveness of character-level similarity metrics may differ for non-Latin scripts or morphologically complex languages. PA-Tool incurs one-time computational overhead from candidate generation and evaluation. While negligible compared to fine-tuning, this cost could become significant for tool libraries with thousands of components.
Appendix A Detailed Algorithm for PA-Tool
Algorithm 1 Pseudocode for PA-Tool
1:Language model $\mathcal{M}$, Component description $d$, samples $s$, temperature $t$, hyperparameter $\alpha$
2:Representative schema $s^{*}$
3:// Stage 1: Candidate Generation
[…]
6:for $i=1$ to $s$ do
7:  $s_{i}\leftarrow$ Generate schema from $\mathcal{M}$ with temperature $t$
8:  Add $s_{i}$ to $\mathcal{C}$
9:end for
10:// Stage 2: Peakedness Computation
[…]
13:for each $s_{i}$ in $\mathcal{C}$ do
14:  $\phi(s_{i})\leftarrow$ count of schemas in $\mathcal{C}$ with edit distance $\leq\tau$ from $s_{i}$ $\triangleright$ Eq. (2)
15:end for
16:// Stage 3: Schema Selection
[…]
18:if $|\mathcal{C}^{*}$| == 1 then
19:  $s^{*}\leftarrow$ the unique schema in $\mathcal{C}^{*}$
20:else
21:  $s^{*}\leftarrow$ schema in $\mathcal{C}^{*}$ closest to $s_{\text{ref}}$ $\triangleright$ Eq. (4)
22:end if
23:return $s^{*}$
Appendix B Prompt Templates
Candidate Name Generation Prompt Generate a {{ component }} name from the description below. The {{ component }} will be used in a tool agent scenario. Description: {{ description }} If component == "tool": Example: Description: A tool that manages files and directories on the system. Output: file_manager Generate only the name without additional explanation. Elif component == "parameter": Example: Context: Tool: file_manager - A tool for managing files and directories Output: file_path Context: Tool: {{ tool_name }} - {{ tool_description }} Generate only the name without additional explanation.
Figure 5: Component name generation prompt for Stage(1) in Figure 2.
Error Classification Prompt You are an expert in analyzing tool selection errors in language models. Given a query, the ground truth tool(s), the model’s selected tool, and the available tool list, classify the error into ONE of these three categories: 1. Schema Misalignment Error: The model selected a tool name that seems plausible but doesn’t exist in the tool list, OR the model selected a tool with a similar name pattern but different from the correct tool. This indicates the model is following its pretrained naming patterns rather than the actual schema. 2. Functional Confusion Error: The model selected an existing tool from the list that has similar functionality to the correct tool, but is not the right choice. The model understands the query but confuses functionally similar tools. 3. Context Understanding Error: The model failed to understand the query’s intent or context, selecting a tool that is functionally unrelated to what the query requires. Query: {query} Ground Truth Tool: {ground_truth} Model Selected: {model_output} Available Tools: {tool_list} IMPORTANT: Respond with ONLY ONE of these exact phrases: Schema Misalignment Error Functional Confusion Error Context Understanding Error
Figure 6: Error classification prompt for analyzing tool selection errors.
Appendix C Example of Schema Generation Process
We provide a concrete example of how PA-Tool generates pretrain-aligned schemas. Given the original tool name DietTool with description "A tool that simplifies calorie counting, tracks diet, and provides insights from many restaurants and grocery stores…", PA-Tool generates 32 candidate names by sampling at temperature 0.4.
The top candidates by frequency are:
  • diet_tracker: 5 occurrences
  • diet_insights: 4 occurrences
  • calorie_tracker: 3 occurrences
  • nutri_guide: 3 occurrences
  • eatwise: 3 occurrences
  • Others: nutrify, nutri_navigator, diet_planner, etc. (1-2 occurrences each)
Importantly, PA-Tool does not simply select the most frequent candidate. Instead, it computes peakedness by measuring how many similar candidates cluster around each option using edit distance. In this case, PA-Tool selects diet_insights (peakedness=4) rather than the most frequent diet_tracker (5 occurrences), as the former has a tighter cluster of similar variants indicating stronger distributional concentration.
It is also worth noting that greedy decoding (temperature 0) produces nutri_guide, which differs from both the most frequent candidate and PA-Tool’s selection. This illustrates three distinct outcomes: (1) PA-Tool’s peakedness-based selection (diet_insights), (2) the most frequent candidate (diet_tracker), and (3) greedy decoding’s output (nutri_guide). These differences highlight that PA-Tool’s selection mechanism considers distributional concentration rather than simple frequency or single-sample generation.
Appendix D Experimental Setup of Supervised Fine-tuning
Setup. As MetaTool does not include a predefined training split, we construct train/validation/test sets by randomly sampling 60%/20%/20% of all task instances. We fine-tune Llama3.1-8B using the hyperparameters detailed in Table 3 and compare four configurations: (1) Vanilla (Base with original schemas); (2) PA-Tool only; (3) SFT only; (4) SFT + PA-Tool.
Table 3: Training hyperparameters for Llama3.1-8B fine-tuning.
Hyperparameter Value
Base Model Llama3.1-8B
LoRA Rank 32
LoRA Alpha 64
LoRA Dropout 0.05
Learning Rate 5e-5
Epochs 5
GPU NVIDIA A6000 D6 48GB
Appendix E Error Analysis Details
Error taxonomy. We categorize tool selection errors into three types based on their underlying failure modes: (1) Schema Misalignment Error occurs when the model generates a plausible but non-existent tool name, or selects a tool with similar naming patterns but different from the correct one. This indicates the model follows its pretrained conventions rather than the provided schema. For example, when the correct tool is get_user_id, the model might generate get_customer_id—a naming pattern it encountered frequently during pretraining but absent from the current schema. (2) Functional Confusion Error occurs when the model selects an existing tool with similar functionality to the correct tool but is not the right choice. This suggests the model understands the query but confuses functionally similar tools. For instance, when asked to send an email notification, the model might select send_sms instead of the correct send_email, both of which are communication tools but use different channels. (3) Context Understanding Error occurs when the model fails to comprehend the query’s intent, selecting a functionally unrelated tool. For example, when asked to delete a user account, the model might incorrectly select create_user or list_products, indicating a fundamental misunderstanding of the task.
Experimental setup. We implement an automated error analyzer using GPT-4.1-mini OpenAI (2024) to classify errors from Llama3.1-8B in both Base and PA-Tool configurations. The analysis is conducted on MetaTool’s three tool selection subtasks: Similar, Scenario, and Multi-tool. We exclude the Reliability subtask as it specifically tests the "no suitable tool" scenario rather than tool selection errors. The classification prompt is provided in Figure 6.
Appendix F Hyperparameter Detailed Results
Table 4: Effect of Number of Candidates on Model Performance
Model N MetaTool
Similar Scenario Reliab. Multi.
Qwen2.5-3B 16 50.5 59.1 85.7 66.4
32 50.0 58.8 86.2 72.6
64 50.9 59.1 86.1 69.6
128 50.5 58.6 86.0 70.6
256 50.2 59.1 83.2 72.4
Qwen2.5-7B 16 63.2 74.9 86.8 84.3
32 64.1 78.4 88.2 84.9
64 62.9 77.7 88.2 87.7
128 62.3 77.4 85.4 84.1
256 63.0 78.4 81.0 87.1
Llama3.2-3B 16 64.2 65.9 60.4 80.9
32 65.7 67.7 60.6 80.5
64 64.3 65.2 57.5 79.5
128 63.7 65.2 52.5 79.3
256 64.8 65.5 54.2 80.7
Llama3.1-8B 16 68.9 79.5 66.9 88.9
32 70.3 79.9 66.0 88.3
64 68.7 77.8 64.5 87.5
128 67.9 79.4 65.1 88.9
256 67.6 79.8 63.9 88.3
Table 5: Effect of Similarity Threshold on Model Performance
Model $\alpha$ MetaTool
Similar Scenario Reliab. Multi.
Qwen2.5-3B 0.1 50.5 58.3 84.1 72.0
0.2 50.0 58.8 86.2 72.6
0.3 51.0 58.6 78.9 60.2
Qwen2.5-7B 0.1 62.7 76.4 84.8 84.9
0.2 64.1 78.4 88.2 84.9
0.3 63.2 75.8 84.7 82.3
Llama3.2-3B 0.1 62.8 66.3 54.8 68.2
0.2 65.7 67.7 60.6 80.5
0.3 65.0 65.5 59.0 79.9
Llama3.1-8B 0.1 68.3 76.9 65.0 88.1
0.2 70.3 79.9 66.0 88.3
0.3 68.1 78.7 64.4 88.7
Table 6: Effect of Sampling Temperature on Model Performance
Model Temp. MetaTool
Similar Scenario Reliab. Multi.
Qwen2.5-3B 0.2 50.2 58.5 85.0 71.0
0.4 50.0 58.8 86.2 72.6
0.6 51.5 58.9 85.5 70.2
0.8 50.5 59.7 83.9 72.0
1.0 51.7 60.2 85.1 67.2
Qwen2.5-7B 0.2 62.2 75.3 88.8 86.9
0.4 64.1 78.4 88.2 84.9
0.6 62.3 75.8 88.6 86.9
0.8 61.4 77.7 87.5 86.1
1.0 63.1 77.4 90.2 86.7
Llama3.2-3B 0.2 64.1 68.3 55.6 74.2
0.4 65.7 67.7 60.6 80.5
0.6 61.3 66.4 64.6 79.9
0.8 63.5 68.7 54.1 80.3
1.0 64.0 66.3 61.5 77.9
Llama3.1-8B 0.2 69.5 79.5 63.5 88.5
0.4 70.3 79.9 66.0 88.3
0.6 68.4 79.8 62.3 80.9
0.8 67.8 79.4 65.2 88.7
1.0 69.5 76.9 65.6 88.5