### v2

# Adapt Tool Schemas to the Models

Abstract
Small language models (SLMs) enable scalable multi-agent tool systems where multiple SLMs handle subtasks orchestrated by a powerful coordinator. However, they struggle with tool-use tasks, particularly in selecting appropriate tools and identifying correct parameters. A common failure mode is schema misalignment: models hallucinate plausible but nonexistent tool names that reflect naming conventions internalized during pretraining but absent from the provided tool schema. Rather than forcing models to adapt to arbitrary schemas, we propose adapting schemas to align with models’ pretrained knowledge. We introduce PA-Tool (Pretraining-Aligned Tool Schema Generation), a training-free method that leverages peakedness, a signal from contamination detection indicating pretraining familiarity, to rename tool components. By generating multiple candidates and selecting those with the highest peakedness across samples, PA-Tool identifies pretraining-aligned naming patterns. Experiments on MetaTool and RoTBench show improvements of up to 17%, with schema misalignment errors reduced by 80%. PA-Tool enables small models to approach state-of-the-art performance while maintaining computational efficiency in adapting to new tools without retraining. Our work demonstrates that schema-level interventions can unlock the tool-use potential of resource-efficient models by adapting schemas to models rather than models to schemas.  [^1] We will release our code upon publication of the paper.
Don’t Adapt Small Language Models for Tools;
Adapt Tool Schemas to the Models
Jonggeun Lee, Woojung Song, Jongwook Han, Haesung Pyun, Yohan Jo Graduate School of Data Science, Seoul National University {jonggeun.lee, opusdeisong, johnhan00, haesung.pyun, yohan.jo}@snu.ac.kr
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
Tool-augmented language models have become essential components of modern AI systems Qu et al. (2025b). As these systems mature, there is growing interest in deploying small language models (SLMs, typically <<10B), driven by both multi-agent systems where a powerful model orchestrates reasoning while multiple SLMs handle subtasks Belcak et al. (2025) and edge deployment scenarios with strict resource constraints Chen et al. (2025b). In these systems, SLMs must perform two critical operations: tool selection (identifying which API to call) and parameter identification (providing appropriate arguments). However, SLMs struggle significantly with these tasks, exhibiting severe performance degradation compared to larger models Patil et al. (2025); Erdogan et al. (2024).
Figure 1: Effect of schema alignment on tool invocation. Top: Models learn tool schema during pretraining. Bottom-Left: When schemas misalign with these internalized patterns, models generate plausible but nonexistent tools. Bottom-Right: Schema alignment with pretrained knowledge prevents such errors.
A common failure pattern is schema misalignment: even when appropriate tools are provided in context, models fail to invoke them and instead hallucinate plausible but nonexistent components (Figure 1, Bottom-Left) within the tool schema—a hierarchical structure of tools and parameters with descriptions. This suggests models are confusing unfamiliar tool schemas with familiar naming conventions internalized during pretraining. Therefore, we hypothesize that aligning tool schemas with patterns from the model’s internalized knowledge can reduce errors and improve tool-calling accuracy (Figure 1, Bottom-Right).
To operationalize this idea, we propose PA-Tool (Pretraining-Aligned Tool Schema Generation), which generates a mapping between original tool component names and model-familiar alternatives. To identify what naming conventions the model is familiar with, we leverage insights from contamination detection methods Dong et al. (2024), which detect whether models encountered specific data during pretraining. A key finding from these studies is that patterns frequently seen during training exhibit peakedness: models generate highly similar outputs across multiple sampling attempts, creating a concentrated distribution Shi et al. (2024a); Li (2023). We adopt this peakedness as a signal for the model’s familiarity with tool names. Concretely, PA-Tool renames each component through three stages: (1) instructing the target language model to generate multiple candidate names based on a description of the component Chen et al. (2021), (2) computing the peakedness of each candidate, measured as the number of other candidates that are sufficiently similar based on character-level edit distance Levenshtein (1965), and (3) selecting the candidate with the highest peakedness as the new name, as it is considered to best align with the model’s pretrained knowledge.
Experiments on MetaTool and RoTBench demonstrate PA-Tool’s effectiveness across diverse scenarios. On MetaTool, PA-Tool achieves improvements up to 17% when models must identify that no suitable tool exists and 9.6% for multi-tool selection scenarios. On RoTBench, performance gains of 5-10% (single-turn) and 4-10% (multi-turn) show that alignment benefits persist across extended contexts. Notably, PA-Tool enables SLMs like Qwen2.5-7B-Instruct to match or exceed Claude Sonnet 4.5 on tool selection tasks.
Our comprehensive analysis reveals that PA-Tool’s effectiveness stems from directly addressing the predominant failure mode in SLMs. Schema misalignment errors, where models generate plausible but nonexistent tool names, decrease by 80.0% with PA-Tool, while other error types show more modest reductions (18.8-24.0%). This targeted impact demonstrates that aligning schemas with pretrained knowledge resolves a bottleneck that disproportionately affects small models. We further validate that peakedness reliably captures learning exposure by showing it consistently increases as models encounter tool schemas during training. PA-Tool also demonstrates broad compatibility with existing approaches: it provides complementary gains when combined with retrieval-based correction, constrained generation, and description enhancement methods, and remains effective even after supervised fine-tuning. Notably, PA-Tool alone outperforms fine-tuned models on several subtasks without any training, while fine-tuning risks catastrophic forgetting to specific tool distributions. Beyond tool selection benchmarks, PA-Tool improves end-to-end task completion on $\tau$-Bench, confirming that schema alignment translates to practical performance gains.
This training-free approach requires only a one-time schema mapping: rather than forcing models to conform to arbitrary documentation, we adapt the interface to match their internalized knowledge, enabling diverse tool use without model modification, retraining, or catastrophic forgetting.
Our contributions are summarized as follows:
  • We propose PA-Tool, a training-free schema optimization method that identifies pretraining-aligned tool names by repurposing peakedness, a signal from contamination detection.
  • We demonstrate improvements of up to 17% across SLMs on MetaTool and RoTBench, with benefits extending from tool selection to parameter identification in both single-turn and multi-turn settings.
  • We demonstrate that through a simple, easily deployable mapping interface, small models can approach closed-source model performance in specific aspects while maintaining computational efficiency.
2 Related Works
2.1 Emerging Agentic Frameworks
Recent advances in LLMs have enabled sophisticated agentic frameworks that decompose complex tasks into specialized modules, each coordinated by dedicated agents Shinn et al. (2023); Madaan et al. (2023); Wu et al. (2025); Agashe et al. (2025); Sapkota et al. (2026). In these multi-agent systems, SLMs are increasingly replacing larger models within individual modules to reduce computational costs and latency while maintaining specialized functionalities Belcak et al. (2025); Cheng et al. (2024). However, these systems remain vulnerable to cascading failures when SLM agents malfunction in foundational tool interaction tasks such as tool selection and parameter identification Erdogan et al. (2024); Patil et al. (2025). A particularly challenging failure mode is schema misalignment: SLM errors often manifest as plausible yet incorrect outputs, such as generating tool names that seem reasonable but don’t exist in the actual schema, making them difficult to detect and correct through traditional validation mechanisms. Our work addresses this vulnerability by generating pretraining-aligned schemas that improve SLM reliability in these critical operations.
2.2 Improving Tool Utilization in LLMs
As LLMs demonstrate the capability to interact with external tools Schick et al. (2023); Hsieh et al. (2023), research has focused on evaluating and improving their tool-use capabilities. Evaluation efforts have developed benchmarks ranging from fine-grained assessments of tool selection and parameter identification Huang et al. (2024); Li et al. (2023a); Ye et al. (2024); Chen et al. (2024); Patil et al. (2025) to end-to-end multi-step evaluation Seo et al. (2025); Yao et al. (2025); Trivedi et al. (2024); Qin et al. (2024); Shim et al. (2025).
Approaches to improving tool-use capabilities follow two main directions. Training-based methods employ supervised fine-tuning Qin et al. (2024); Liu et al. (2025); Zhang et al. (2025) or reinforcement learning Qian et al. (2025); Zhou et al. (2025); Shi et al. (2024b); Feng et al. (2025); Chen et al. (2025a), but they require substantial data or computational resources. Training-free methods refine tool documentation Yuan et al. (2025); Qu et al. (2025a) or leverage interaction histories Wang et al. (2024); Cui et al. (2025); Zhao et al. (2024); Fu et al. (2024). However, existing training-free approaches primarily focus on improving descriptions rather than aligning the schema itself with model-preferred representations. Our work directly addresses this gap by generating schemas aligned with pretraining distributions, targeting the root cause of schema misalignments.
2.3 Data Contamination Detection
Data contamination, defined as overlap between pretraining data and evaluation benchmarks, can inflate performance through memorization. Early detection methods use n-gram overlap Brown et al. (2020), but these approaches fail to detect semantic paraphrasing. Probability-based approaches such as Min-k% Prob Shi et al. (2024a) and perplexity analysis Li (2023) require token probabilities that are unavailable in closed-source models. LLM Decontaminator Yang et al. (2023) uses auxiliary models for semantic similarity but introduces additional dependencies. Most relevant to our work, CDD Dong et al. (2024) identifies memorized patterns by measuring peakedness in sampled candidates, making it applicable to any black-box model. We adapt this peakedness mechanism to identify pretraining-aligned schema patterns, transforming contamination detection into a constructive tool for schema optimization.
Figure 2: Overview of our PA-Tool framework.
3 PA-Tool : Pretraining-Aligned Tool Schema Generation
We now formalize our approach to generating pretraining-aligned tool schemas. A tool schema is structured documentation that defines tools and parameters along with descriptions for each component. Our objective is to rename these components with names that align with patterns the model has frequently encountered during pretraining.
3.1 Framework Overview
Let $\mathcal{M}$ denote a pre-trained language model. Given a natural language description $d$ of a schema component’s functionality (e.g., a tool or a parameter description), our objective is to generate an optimal name $s^{*}$ that represents the naming pattern most deeply internalized by $\mathcal{M}$. As illustrated in Figure 2, our approach operates in three stages: (1) we provide $\mathcal{M}$ with a description of a component and instruct it to generate its name multiple times (Stage 1), (2) we compute peakedness scores that measure how many similar candidates cluster around each candidate name (Stage 2), and (3) we select the candidate with the highest peakedness as the pretraining-aligned name (Stage 3). By applying this process to each component in the schema hierarchy, we construct a dictionary mapping from original names to pretraining-aligned names. The complete algorithm is provided in Appendix A.
3.2 Stage 1: Candidate Generation
In this stage, we collect diverse candidate names of a component that the model may have encountered during pretraining. As illustrated in Figure 2-1, given a component’s description, we sample $N$ candidate names $\mathcal{C}=\{s_{1},s_{2},\ldots,s_{N}\}$ from the language model with temperature $t\in(0,1]$. This temperature-controlled sampling explores the model’s learned distribution beyond the single greedy path, revealing diverse naming patterns Chen et al. (2021). The detailed prompt structure is provided in Appendix B. Additionally, we generate a reference name $s_{\text{ref}}$ using greedy decoding ($t=0$) for tie-breaking purposes in the selection stage.
3.3 Stage 2: Peakedness Computation
Following contamination detection principles Dong et al. (2024), we analyze the local concentration of the output distribution to identify strongly memorized patterns. As illustrated in Figure 2-2, for each candidate name $s_{i}$, we compute its peakedness score by counting how many similar candidate names the model generates around it. The intuition is that candidate names with many similar variants indicate pretrain-frequent patterns that the model generates consistently, while isolated candidates suggest less familiar patterns.
To quantify this clustering behavior, we first define a similarity threshold $\tau$ based on the maximum character length in the candidate set:
$\tau=\alpha\cdot\ell_{\max}$ (1)
where $\ell_{\max}=\max_{s_{i}\in\mathcal{C}}|s_{i}|$ is the maximum character length across all candidates, and $\alpha\in[0,1]$ is a hyperparameter that controls the strictness of similarity. This length-adaptive threshold ensures that longer names are allowed proportionally more variation while maintaining consistent similarity criteria across different name lengths.
The peakedness score for each candidate $s_{i}$ is then computed as:
$\phi(s_{i})=\sum_{j\neq i}\mathbb{I}(d_{\text{edit}}(s_{i},s_{j})\leq\tau)$ (2)
where $d_{\text{edit}}(\cdot,\cdot)$ denotes the character-level edit distance using the Levenshtein algorithm Levenshtein (1965), and $\mathbb{I}(\cdot)$ is the indicator function that equals 1 when the condition is satisfied and 0 otherwise. This score counts the number of candidates that fall within the similarity threshold from $s_{i}$, reflecting how many similar names the model generates around this pattern.
3.4 Stage 3: Representative Name Selection
The representative name is selected as the candidate with the maximum peakedness:
$s^{*}=\underset{s_{i}\in\mathcal{C}}{\mathrm{arg\,max}}\;\phi(s_{i})$ (3)
This criterion identifies the naming pattern that the model generates most consistently across multiple sampling attempts, indicating it represents the most deeply internalized pattern from the training data. In case of ties where multiple candidates achieve the same maximum peakedness score, we break ties by selecting the candidate with the minimum edit distance to the reference name:
$s^{*}=\underset{s_{i}\in\mathcal{C}^{*}}{\mathrm{arg\,min}}\;d_{\text{edit}}(s_{i},s_{\text{ref}})$ (4)
where $\mathcal{C}^{*}=\{s_{i}\in\mathcal{C}:\phi(s_{i})=\max_{s_{j}\in\mathcal{C}}\phi(s_{j})\}$ contains all candidates with maximum peakedness.
This approach rests on the hypothesis that frequently occurring patterns in training data create local maxima in the model’s output distribution. By identifying regions of high peakedness, we effectively locate these memorized naming conventions, which represent the most natural and well-formed component names according to the model’s learned knowledge. Through iterative application across all schema components (Figure 2-3), we obtain the final pretraining-aligned schema. An example of a schema generated through this process is provided in Appendix C. We repeat our three-stage process for all tool sets within the system. When name collisions occur across different tools due to highly similar descriptions, we resolve them using the priority-based locking mechanism (Appendix I).
4 Experimental Setup
Model Method MetaTool RoTBench
Tool Selection Single-turn Multi-turn
Similar Scenario Reliability Multi-tool Tool Sel. Param Iden. Tool Sel. Param Iden.
Closed-Source Models
GPT-4.1-mini Base 79.6 84.3 76.3 72.2 79.1 58.1 28.6 12.9
Gemini-2.5-Flash Base 70.0 79.8 89.2 77.3 82.9 56.2 22.9 12.9
Claude-Sonnet-4.5 Base 75.6 83.0 84.3 85.1 83.3 67.6 25.7 17.1
Open-Source Models
Qwen2.5-3B Base 48.7 55.3 83.6 75.1 1.9 0.0 1.4 0.0
Greedy 49.1 57.0 82.9 63.4 6.7 3.8 5.7 1.4
MostFreq 50.5 56.3 84.3 73.0 12.4 9.5 5.7 2.9
Human 54.6 62.9 80.6 72.6 - - - -
PA-Tool 50 58.8 86.2 72.6 6.7 1.9 7.1 4.3
Qwen2.5-7B Base 59.6 74.4 78.3 78.3 49.5 20 8.6 5.7
Greedy 60.6 75.4 85.4 82.7 50.5 21 14.3 8.6
MostFreq 65.3 75.5 90.0 86.1 50.5 23.8 12.9 10
Human 65.8 76.7 78.6 82.1 - - - -
PA-Tool 64.1 78.4 88.2 84.9 55.2 21.9 35.7 11.4
Llama3.2-3B Base 55.0 58.6 43.6 79.1 56.2 20.0 27.1 5.7
Greedy 57.7 58.9 39.8 70.4 59.1 20.0 28.6 4.3
MostFreq 61.3 64.2 65.7 81.3 60.0 18.1 32.9 7.1
Human 58.6 60.9 55.2 72.6 - - - -
PA-Tool 65.7 67.7 60.6 80.5 62.9 21.9 31.4 5.7
Llama3.1-8B Base 61.5 73.9 53.5 78.7 58.1 17.1 17.1 4.3
Greedy 64.6 72.9 51.5 78.9 63.8 18.1 17.1 4.3
MostFreq 68.8 79.3 66.4 85.7 66.7 18.1 27.1 4.3
Human 69.3 78.9 63.5 86.5 - - - -
PA-Tool 70.4 79.9 66.0 88.3 68.6 18.1 22.9 4.3
Table 1: Performance comparison on MetaTool and RoTBench. All metrics are reported as accuracy (%). Bold indicates the best performance within each open-source model and among all closed-source models.
4.1 Benchmarks
MetaTool. MetaTool (Huang et al., 2024) evaluates tool selection capabilities across 4,287 test cases in four subtasks: (1) Similar tests semantic comprehension by distinguishing tools with overlapping functionalities (e.g., Sudoku vs. Tic-Tac-Toe); (2) Scenario selects appropriate tools based on user-specific contexts and requirements (e.g., software engineers, students); (3) Reliability assesses whether models can directly indicate when no suitable tool is available rather than hallucinating; and (4) Multi-tool measures whether models can correctly select multiple tools when tasks require composition of functionalities.
RoTBench. RoTBench (Ye et al., 2024) evaluates whether models can robustly use tools when descriptions contain various noise perturbations (e.g. reversed names). However, we focus exclusively on the Clean environment to isolate the effects of PA-Tool on core tool-use capabilities rather than noise robustness. We evaluate on 105 test cases across 568 tools in both single-turn and multi-turn settings. RoTBench measures two fundamental tool use capabilities: (1) Tool Selection evaluates whether models correctly identify the appropriate tool; and (2) Parameter Identification assesses whether models accurately extract the required parameter set, conditioned on correct tool selection.
4.2 Models and Baselines
We focus on models with 8B parameters or fewer (SLMs), evaluating four open-source models: Qwen 2.5 3B/7B (Qwen et al., 2025), Llama-3.1 8B (Meta AI, 2024a), and Llama 3.2 3B (Meta AI, 2024b). For each model, we evaluate five configurations: (1) Base uses the original schema without modifications; (2) Greedy uses greedy decoding to generate deterministic schema names (no schema-change control); (3) MostFreq selects the most frequently generated candidate from our sampling process; (4) Human uses schemas where two PhD-level researchers with software engineering backgrounds manually renamed tools to more intuitive names (evaluated only on MetaTool due to RoTBench’s large tool-sets); (5) PA-Tool applies our proposed peakedness-based method. To contextualize SLM performance and quantify PA-Tool’s improvements, we also compare against three state-of-the-art closed-source models: GPT-4.1-mini (OpenAI, 2024), Gemini 2.5 Flash (Basu Mallick et al., 2025), and Claude Sonnet 4.5 (Anthropic, 2025).
4.3 Implementation Details
When generating schemas with PA-Tool, we use 32 candidates at temperature 0.4 with $\alpha=0.2$ (sensitivity analysis in Appendix J). Each component’s description is provided individually (See Figure 2-1; prompt template in Appendix B).
For benchmark inference, we use temperature 0 to ensure reproducible results across all experiments. We use accuracy as the primary evaluation metric across both benchmarks, measuring the percentage of test cases where the model’s predictions exactly match the ground-truth labels. Detailed evaluation protocols are in Appendix D.
5 Main Results
Table 1 presents comprehensive results across MetaTool and RoTBench benchmarks, demonstrating PA-Tool’s effectiveness.
MetaTool.
PA-Tool substantially improves performance over Base models across all four MetaTool subtasks. The most substantial gains appear in Reliability (up to 17.0%, e.g., Llama3.2-3B: 43.6→60.6%), where models must recognize when no suitable tool exists, a task critically dependent on clearly understanding available options. In Multi-tool, gains reach 9.6% (Llama3.1-8B: 78.7→88.3%). When tasks require identifying multiple tools simultaneously, schema misalignment compounds across each selection, making alignment particularly critical. Similar and Scenario tasks show improvements up to 10.7%.
PA-Tool consistently outperforms training-free alternatives. Greedy occasionally underperforms Base (Llama3.2-3B Reliability: 39.8% vs. 43.6%) by generating only a single candidate, limiting schema space exploration. While MostFreq sometimes achieves competitive results by capturing frequency, PA-Tool shows more consistent improvements by measuring distributional concentration. Compared to Human-designed schemas, PA-Tool achieves comparable or superior performance (Llama3.2-3B Similar: 65.7% vs. 58.6%), demonstrating automated alignment can match human intuition while being scalable.
RoTBench.
RoTBench evaluates models in single-turn and multi-turn settings across tool selection and parameter identification. PA-Tool demonstrates consistent improvements of 5-10% in single-turn tool selection, with gains of up to 10% remaining in multi-turn settings, demonstrating effectiveness across extended conversations.
MostFreq shows strong performance for some models in single-turn (Qwen2.5-3B: 12.4%), but PA-Tool demonstrates more consistent improvements in multi-turn settings and across larger models, suggesting better generalization.
Comparison with Closed-source Models.
While a performance gap remains versus state-of-the-art closed-source models (Claude Sonnet 4.5: 83.3% tool selection on RoTBench single-turn), PA-Tool enables small models to achieve competitive or superior performance in specific aspects. In MetaTool’s Multi-tool, Llama3.1-8B with PA-Tool (88.3%) surpasses all closed-source models including Claude Sonnet 4.5 (85.1%). In Reliability, Qwen2.5-7B with PA-Tool (88.2%) approaches Gemini 2.5 Flash (89.2%). These results demonstrate that targeted schema alignment can enable resource-efficient models to match state-of-the-art systems where naming conventions critically impact performance.
Generalization to Diverse Models.
Beyond the models examined here, we conduct additional experiments on diverse model families and scales, including both SLMs (Ministral-8B, GPT-4.1-nano) and LLMs (Llama-3.3-70B, GPT-4.1-mini, Gemini-2.5-Flash). PA-Tool shows substantial improvements for SLMs with consistent gains up to 11.2% on RoTBench. For LLMs, improvements are generally more modest, though notable exceptions exist—GPT-4.1-mini achieves 12.0% gain on MetaTool Multi-tool, demonstrating that even highly capable models benefit from schema alignment in certain tasks (see Appendix E for details).
6 Analysis
6.1 Error Analysis
Figure 3: Error type distribution for Llama-3.1-8B on MetaTool tool selection tasks.
To understand how PA-Tool addresses different failure modes, we analyze error distributions in Llama3.1-8B before and after applying PA-Tool. We categorize incorrect predictions into three types: Schema Misalignment Error (generating nonexistent but plausible tools), Functional Confusion Error (selecting wrong tools with similar functionality), and Context Understanding Error (selecting functionally unrelated tools). Detailed definitions and settings are in Appendix G.
Figure 3 shows how PA-Tool affects different error types. PA-Tool substantially reduces all error types, with particularly strong impact on Schema Misalignment (80.0% reduction) and meaningful improvements on Functional Confusion (24.0%) and Context Understanding (18.8%). This demonstrates that schema alignment not only eliminates naming-related failures but also indirectly improves the model’s ability to distinguish tools and understand contexts, likely by reducing cognitive load from misaligned names.
6.2 Integration with Training-Free Methods
Beyond schema renaming, we evaluate PA-Tool’s synergy with other training-free approaches. We test three categories of methods: (1) Retrieval-based correction, which maps misaligned outputs to valid tools post-hoc using BM25 (Robertson and Zaragoza, 2009) or ToolLLM (Qin et al., 2024); (2) Constrained generation, which enforces valid tool names through JSON schema constraints during decoding; and (3) Description enhancement (EasyTool (Yuan et al., 2025)), which rewrites tool descriptions for clarity without modifying names. These methods address different aspects of tool-use and are compatible with PA-Tool’s schema alignment approach. We evaluate both standalone and combined configurations (Method + PA-Tool), with detailed settings presented in Appendix K.
Table 11 shows that retrieval methods show limited gains on their own (<3% improvement). Even with post-hoc correction opportunities, matching often fails to capture the model’s intended tool. Constrained generation substantially improves RoTBench (Qwen2.5-7B: 78.1% vs. Base 49.5%) by eliminating format errors, but shows mixed results on MetaTool.
However, integrating PA-Tool with these methods consistently improves both settings (Table 12), confirming the two approaches are complementary. PA-Tool consistently outperforms EasyTool (Table 13), with notable gaps on MetaTool (e.g., Llama3.2-3B Similar: 65.7% vs. 45.3%). This suggests naming alignment is more effective than description enhancement. Moreover, rewriting descriptions risks altering perceived tool functionality, while PA-Tool preserves semantics by modifying only names. Nevertheless, both remain compatible, with some combined configurations showing competitive results. Across all methods, PA-Tool demonstrates broad compatibility and provides complementary gains when combined, confirming it addresses a distinct failure mode.
6.3 Integration with Supervised Fine-tuning
[…]
Table 2 shows PA-Tool provides consistent gains across all configurations. Notably, PA-Tool alone outperforms both SFT1 and SFT2 on Reliability and Multi-tool without any training. SFT2 shows marginal gains over SFT1 on MetaTool but degrades on RoTBench (59.1% vs. 61.0%), indicating catastrophic forgetting. Combined configurations achieve the best results, with SFT2 + PA-Tool reaching 73.2% on Similar. These results reveal two key insights. First, pretraining-aligned naming preferences persist through fine-tuning, as schemas generated from base or fine-tuned models are similarly effective. Second, SFT’s cross-domain degradation highlights its limitation: it specializes to specific tool distributions, risking catastrophic forgetting. PA-Tool’s training-free nature avoids this, providing robust improvements across both in-domain and out-of-domain benchmarks.
Configuration MetaTool RoTBench
Similar Scenario Reliab. Multi. Tool Sel.
Base 65.2 71.1 55.1 79.8 58.1
+ PA-Tool 72.7 75.9 67.2 89.9 68.6
+ SFT1 71.2 77.6 57.1 82.8 61.0
+ SFT1 + PA-Tool 72.7 80.8 57.1 89.9 65.7
[…]
+ SFT2 + PA-Tool 73.2 79.2 55.6 86.9 61.9
Table 2: Results of fine-tuning and PA-Tool on 20% random sample from MetaTool and full RoTBench.
Figure 4: Peakedness across training epochs.
6.4 Validating Peakedness
PA-Tool assumes that peakedness reflects models’ familiarity with naming patterns. To validate this, we simulate pretraining conditions where models encounter tool schemas and measure whether peakedness increases across training epochs. We use the same MetaTool training data and experimental settings from Section 6.3 but with a different loss computation: unlike SFT which computes loss only on assistant responses, we compute loss on all tokens to simulate pretraining conditions. We train four models (Qwen2.5-3B/7B, Llama3.2-3B, Llama3.1-8B), measuring peakedness at epochs 0, 10, 20, 30, 50, and 100 with 256 candidates per tool component for stable measurements. Figure 4 shows that peakedness consistently increases with training epochs across all models, with gains up to +25.8% (Llama3.2-3B). This validates our hypothesis that peakedness captures learning exposure, supporting PA-Tool’s use of peakedness to identify pretraining-aligned naming conventions.
6.5 Additional Benchmarks
To demonstrate that PA-Tool generalizes across different evaluation settings, we evaluate on API-Bank Li et al. (2023b) and $\tau$-Bench Yao et al. (2025). API-Bank evaluates tool-calling accuracy by measuring exact matches between predicted and ground-truth tools and parameters. $\tau$-Bench is an end-to-end benchmark that assesses agent performance through multi-turn dialogues where agents must complete tasks through interaction with users.
API-Bank. PA-Tool consistently improves performance in most settings (Table 3). Qwen2.5-3B shows substantial gains on the Call task (18.0→28.5), while improvements persist on Call+Retrieve, which requires both tool selection and information retrieval. This confirms that schema alignment enhances tool-calling accuracy.
$\tau$-Bench. To verify that improved tool selection translates to better task completion, we evaluate on $\tau$-Bench (Retail). We run each evaluation 5 times to mitigate variance from the user simulator. PA-Tool consistently improves end-to-end task completion across all models (Table 3). Analyzing Llama3.1-8B’s tool-calling trajectories reveals that schema misalignment errors decreased from 115 to 98 cases, demonstrating that addressing schema misalignment contributes substantially to the observed performance gains.
Call Call+Retrieve $\tau$-Bench
Model Base PA-Tool Base PA-Tool Base PA-Tool
Qwen2.5-3B 18.0 28.5 18.5 21.9 3.7 3.8
Qwen2.5-7B 25.7 34.7 23.5 26.9 6.8 9.7
Llama3.2-3B 5.1 5.4 10.9 9.2 4.5 5.6
Llama3.1-8B 28.0 29.8 21.0 22.4 9.7 11.1
Table 3: Performance on API-Bank and end-to-end task completion (%) on $\tau$-Bench (Retail, N=5).
7 Conclusion
We introduced PA-Tool, a training-free method that aligns tool schemas with models’ pretrained knowledge by leveraging peakedness as a pretraining familiarity signal. Experiments demonstrate improvements of up to 17% with schema misalignment errors reduced by 80%, validating schema adaptation as an effective strategy for enhancing tool use in small language models. PA-Tool’s practical advantages make it particularly valuable for resource-constrained deployments—as a simple schema-level intervention, it can be applied without model training or fine-tuning, requiring only straightforward name mapping. By bridging pretrained knowledge and real-world tool interfaces, PA-Tool unlocks small models’ potential for tool-augmented applications while preserving computational efficiency.
Limitations
The effectiveness of PA-Tool varies with model capacity in an interpretable way: smaller models show consistent improvements, as schema misalignment remains a primary failure mode. In contrast, larger models exhibit diminishing gains as parameter count increases, likely due to their stronger reasoning abilities. However, performance improvements remain model-dependent; GPT-4.1-mini still achieves a 12.0% improvement on the Multi-tool task in Metatool, suggesting that certain architectures benefit from schema alignment regardless of scale.
Our reliance on peakedness as an alignment signal assumes this metric reliably indicates pretraining familiarity. While validated across our experiments, this relationship may vary for models with substantially different training distributions. Additionally, our evaluation focuses on English-language schemas; the effectiveness of character-level similarity metrics may differ for non-Latin scripts or morphologically complex languages. PA-Tool incurs one-time computational overhead from candidate generation and evaluation. While negligible compared to fine-tuning, this cost could become non-trivial for tool libraries with thousands of components.
Appendix A Detailed Algorithm for PA-Tool
Algorithm 1 Pseudocode for PA-Tool
1:Language model $\mathcal{M}$, Component description $d$, samples $s$, temperature $t$, hyperparameter $\alpha$
2:Representative schema $s^{*}$
3:// Stage 1: Candidate Generation
[…]
6:for $i=1$ to $s$ do
7:  $s_{i}\leftarrow$ Generate schema from $\mathcal{M}$ with temperature $t$
8:  Add $s_{i}$ to $\mathcal{C}$
9:end for
10:// Stage 2: Peakedness Computation
[…]
13:for each $s_{i}$ in $\mathcal{C}$ do
14:  $\phi(s_{i})\leftarrow$ count of schemas in $\mathcal{C}$ with edit distance $\leq\tau$ from $s_{i}$ $\triangleright$ Eq. (2)
15:end for
16:// Stage 3: Schema Selection
[…]
18:if $|\mathcal{C}^{*}$| == 1 then
19:  $s^{*}\leftarrow$ the unique schema in $\mathcal{C}^{*}$
20:else
21:  $s^{*}\leftarrow$ schema in $\mathcal{C}^{*}$ closest to $s_{\text{ref}}$ $\triangleright$ Eq. (4)
22:end if
23:return $s^{*}$
Appendix B Prompt Templates
Candidate Name Generation Prompt Generate a {{ component }} name from the description below. The {{ component }} will be used in a tool agent scenario. Description: {{ description }} If component == "tool": Example: Description: A tool that manages files and directories on the system. Output: file_manager Generate only the name without additional explanation. Elif component == "parameter": Example: Context: Tool: file_manager - A tool for managing files and directories Output: file_path Context: Tool: {{ tool_name }} - {{ tool_description }} Generate only the name without additional explanation.
Figure 5: Component name generation prompt for Stage(1) in Figure 2.
Error Classification Prompt You are an expert in analyzing tool selection errors in language models. Given a query, the ground truth tool(s), the model’s selected tool, and the available tool list, classify the error into ONE of these three categories: 1. Schema Misalignment Error: The model selected a tool name that seems plausible but doesn’t exist in the tool list, OR the model selected a tool with a similar name pattern but different from the correct tool. This indicates the model is following its pretrained naming patterns rather than the actual schema. 2. Functional Confusion Error: The model selected an existing tool from the list that has similar functionality to the correct tool, but is not the right choice. The model understands the query but confuses functionally similar tools. 3. Context Understanding Error: The model failed to understand the query’s intent or context, selecting a tool that is functionally unrelated to what the query requires. Query: {query} Ground Truth Tool: {ground_truth} Model Selected: {model_output} Available Tools: {tool_list} IMPORTANT: Respond with ONLY ONE of these exact phrases: Schema Misalignment Error Functional Confusion Error Context Understanding Error
Figure 6: Error classification prompt for analyzing tool selection errors.
Training Data Generation Prompt You are an expert at explaining tool selection reasoning for AI systems. Your task is to generate high-quality, flowing reasoning that explains WHY specific tools were chosen for a given query. Your Task: Given: A user’s query A list of available tools with descriptions The tools that were selected (answer) Generate: A clear, logical reasoning in natural paragraph form that explains why these specific tools are the best choices. Quality Criteria: Your reasoning MUST: 1. Analyze the query deeply: Identify key requirements, implicit needs, and user intent 2. Connect query to tools: Explicitly link query elements to tool capabilities 3. Justify each selection: Explain why each tool is necessary 4. Show tool synergy: Explain how the tools work together 5. Consider alternatives: Briefly mention why other tools were NOT chosen 6. Be specific: Use concrete details from the query and tool descriptions 7. Write in flowing prose: Use natural paragraphs, not bullet points or lists Example: Now Generate: User’s Query: {{ query }} Selected Tools: {{ selected_tools }} Write your reasoning as flowing paragraphs that naturally cover query analysis, tool justification for each selected tool, their combined strategy, and why alternatives were not chosen.
Figure 7: Training data generation prompt for supervised fine-tuning in Metatool.
Prompt template used for inference on MetaTool You are a helpful AI assistant. Your current task is to choose the appropriate tool to solve the user’s query based on their question. I will provide you with the user’s question and information about the tools. If there is a tool in the list that is applicable to this query, please return the name of the tool (you can only choose one tool). If there isn’t, please return ‘None.’ Additionally, you will need to support your answer with a brief explanation. User’s Query: [User’s Query Start] Planning a beach day next week, while considering the ideal weather conditions, availability of beach accessories, preferred beach location, potential activities to engage in like swimming, sunbathing, or playing beach games, and coordinating with friends or family members to join in the fun. [User’s Query End] List of Tools with Names and Descriptions: [List of Tools with Names and Descriptions Start] 1. tool name: airqualityforeast, tool description: [‘Planning something outdoors? Get the 2-day air quality forecast for any US zip code.’] 2. tool name: WeatherTool, tool description: Provide you with the latest weather information. 3. tool name: AusSurfReport, tool description: [‘Get today’s surf report for any break throughout Australia!’] 10. tool name: EarthquakeTool, tool description: Provides real-time earthquake notifications and news. [List of Tools with Names and Descriptions End] Here are some examples: [Examples Start] query: “I’m planning a hiking trip next week. What will the weather be like in the Grand Canyon?” tool: WeatherTool query: “I’m a beginner surfer. Can you suggest a beach with mild waves for me to practice in Australia?” tool: AusSurfReport [Examples End] User query: “Planning a beach day next week, while considering the ideal weather conditions…” tool:
Figure 8: Prompt template used for model inference on MetaTool benchmark. The prompt includes task instructions, a list of candidate tools with descriptions, few-shot examples, and the target user query.
Prompt template used for model inference on RoTBench System: You are an expert in using tools. At each step, analyze the state and decide the next action with a function call. Available tools: [{"name": "get_translation_nllb", "description": "Translate text using NLLB model.", "parameters": {"input_text": {...}, "tgt_lang": {...}, "src_lang": {...}, ...}}, {"name": "get_translation_baidu", "description": "Translate using BAIDU API.", "parameters": {"text": {...}, "tgt_lang": {...}, ...}}, ...] User: “What is the translation of ‘See you later’ in Japanese?” Expected output: Thought: Need to translate English to Japanese Action: get_translation_nllb Action Input: {"input_text": "See you later", "tgt_lang": "jpn_Jpan", "src_lang": "eng_Latn", …}
Figure 9: Prompt template used for model inference on RoTBench. The prompt includes a system message, available tool schemas with full parameter specifications, and the user query. Models must output structured responses with reasoning, tool selection, and parameter values.
Prompt template used for model inference on API-Bank Instruction: Based on the given API description and the existing conversation history 1..t, please generate the API request that the AI should call in step t+1 and output it in the format of [ApiName(key1=’value1’, key2=’value2’, …)], replace the ApiName with the actual API name, and replace the key and value with the actual parameters. Your output should start with a square bracket “[” and end with a square bracket “]”. Do not output any other explanation or prompt or the result of the API call in your output. This year is 2023. API descriptions: [{"name": "GetWeather", "description": "Get current weather information", "parameters": {"location": {"type": "string", "description": "City name"}, "unit": {"type": "string", "enum": ["celsius", "fahrenheit"]}}}, ...] Input: User: What’s the weather like in San Francisco? AI: I’$l$ check the current weather in San Francisco for you. Expected output: [GetWeather(location=’San Francisco’, unit=’fahrenheit’)]
Figure 10: Prompt template used for model inference on API-Bank. The prompt includes instructions for API call generation based on conversation history, API schemas with parameter specifications, and example dialogue with expected API call output in bracket notation.
Prompt template used for model inference on Tau-Bench Instruction: You need to act as an agent that uses the above tools to help the user according to the above policy. At each step, your generation should have exactly the following format: Thought: <A single line of reasoning to process the context and inform the decision making. Do not include extra lines.> Action: {"name": <The name of the action>, "arguments": <The arguments to the action in json format>} The Action will be parsed, so it must be valid JSON. You should not use made-up or placeholder arguments. Available tools: [{"type": "function", "function": {"name": "get_current_weather", "description": "Get the current weather", "parameters": {"type": "object", "properties": {"location": {"type": "string", "description": "The city and state, e.g. San Francisco, CA"}, "format": {"type": "string", "enum": ["celsius", "fahrenheit"], "description": "The temperature unit to use. Infer this from the users location."}}, "required": ["location", "format"]}}}, ...] User: “I want to know the current weather of San Francisco” Expected output: Thought: Since the user asks for the weather of San Francisco in USA, the unit should be in fahrenheit. I can query get_current_weather to get the weather. Action: {"name": "get_current_weather", "arguments": {"location": "San Francisco, CA", "format": "fahrenheit"}} [If tool returns "70F"] Thought: I can answer the user now. Action: {"name": "respond", "arguments": {"response": "The current weather of San Francisco is 70F."}}
Figure 11: Prompt template used for model inference on Tau-Bench. The prompt includes instructions for structured agent behavior, available tool schemas with parameter specifications, and example user query with expected multi-step reasoning and tool invocations.
Appendix C Example of Schema Generation Process
We provide a concrete example of how PA-Tool generates pretraining-aligned schemas. Given the original tool name DietTool with description "A tool that simplifies calorie counting, tracks diet, and provides insights from many restaurants and grocery stores…", PA-Tool generates 32 candidate names by sampling at temperature 0.4.
The top candidates by frequency are:
  • diet_tracker: 5 occurrences
  • diet_insights: 4 occurrences
  • calorie_tracker: 3 occurrences
  • nutri_guide: 3 occurrences
  • eatwise: 3 occurrences
  • Others: nutrify, nutri_navigator, diet_planner, etc. (1-2 occurrences each)
Importantly, PA-Tool does not simply select the most frequent candidate. Instead, it computes peakedness by measuring how many similar candidates cluster around each option using edit distance. In this case, PA-Tool selects diet_insights (peakedness=4) rather than the most frequent diet_tracker (5 occurrences), as the former has a tighter cluster of similar variants indicating stronger distributional concentration.
It is also worth noting that greedy decoding (temperature 0) produces nutri_guide, which differs from both the most frequent candidate and PA-Tool’s selection. This illustrates three distinct outcomes: (1) PA-Tool’s peakedness-based selection (diet_insights), (2) the most frequent candidate (diet_tracker), and (3) greedy decoding’s output (nutri_guide). These differences highlight that PA-Tool’s selection mechanism considers distributional concentration rather than simple frequency or single-sample generation.
Figure 12: Detailed performance gain (PA-Tool - Base) across different models and metrics for MetaTool dataset.
Figure 13: Detailed performance gain (PA-Tool - Base) across different models and metrics for RoTBench dataset.
Appendix D Evaluation Protocol
D.1 MetaTool
In MetaTool evaluation, the agent receives (1) a user query, (2) a list of candidate tools with their names and descriptions, and (3) few-shot examples. The agent must select the appropriate tool name(s) from the provided list. For the Similar, Scenario, and Multi-tool subtasks, the agent selects one or more ground-truth tools. For the Reliability subtask, when no suitable tool exists, the agent must output “None” to indicate the query cannot be solved with the given tools. Critically, the multi-candidate selection nature of the task is preserved: models must reason about and distinguish between all available tools in the candidate list, regardless of their names.
Prompt Structure.
We present an example prompt template used for model inference (Figure 8).
Scoring Methodology.
We revised the original MetaTool scoring methodology to address several edge cases and improve evaluation accuracy. The original method relied heavily on manual analysis and did not adequately handle ambiguous cases where models produced partially correct outputs or included both correct and incorrect tools in their responses.
For single-tool tasks (similar and scenario tasks), when exactly one tool match is found, we now first verify that it corresponds to the ground-truth label before considering it correct. Additionally, if the keyword “None” appears in the model output, we check its position relative to the matched tool. The response is marked correct only if the matched tool appears before “None,” as this indicates the model prioritized the correct tool over rejection. For cases with two or more matches, we replaced manual analysis with automated evaluation using GPT-4.1-mini, which assesses whether the model output selected only the ground-truth tools. This change was necessary because manual review became impractical at scale and introduced inconsistencies in scoring.
For reliability tasks, we introduced logic to verify correct “None” predictions. If no matches are found and “None” is present, or if matches exist but “None” appears before them, the response is marked correct. This refinement was added because the original methodology could not distinguish between appropriate rejections and cases where models incorrectly included tool names alongside “None.”
For multi-tool tasks, responses with exactly one match are now explicitly marked incorrect, as these tasks require selecting multiple tools. Cases with more than two matches are evaluated using GPT-4.1-mini instead of manual review, ensuring consistent and scalable assessment. These refinements enable more precise automated scoring while capturing nuanced model behaviors that the original methodology could not adequately assess.
D.2 RoTBench
In RoTBench evaluation, the agent receives (1) a system message describing the task format, (2) a list of available tools with complete schemas including names, descriptions, and parameter specifications, and (3) a user query. The agent must provide both tool selection and parameter identification in a structured format.
Prompt Structure.
We present an example prompt template used for model inference (Figure 9).
Evaluation Metrics.
Tool Selection accuracy is measured by whether the model selects the correct tool name from the available list. Parameter Identification accuracy is computed conditionally: only cases where the model correctly selected the tool are evaluated for whether the model also identified the correct set of required parameters. This two-stage evaluation isolates the impact of schema alignment on each capability independently, while preserving the multi-candidate selection challenge where models must reason across multiple tool options.
D.3 API-Bank
API-Bank (Li et al., 2023b) is a comprehensive benchmark for evaluating tool-augmented language models across realistic API usage scenarios. The benchmark implements 73 runnable APIs spanning diverse domains and evaluates models on 314 manually annotated dialogues containing 753 API calls.
API-Bank assesses two fundamental capabilities: (1) Call, where models must correctly invoke APIs with appropriate parameters when APIs are provided in the context, and (2) Retrieval+Call, where models must first retrieve relevant APIs from a large pool using an API Search function, then call the selected APIs. Both tasks require exact matching of tool names and parameters with ground-truth annotations for correctness. We evaluate on both capabilities to assess whether PA-Tool improves tool selection when schemas are either explicitly provided (Call) or must be retrieved (Retrieval+Call). The detailed prompt given to the agent is shown in Appendix 10.
D.4 $\tau$-Bench
$\tau$-Bench (Yao et al., 2025) evaluates agents on realistic, multi-turn task-oriented dialogues where they must interact with simulated users while following domain-specific policies. Unlike traditional benchmarks that provide all information upfront, $\tau$-Bench requires agents to incrementally gather information through conversation, consult policy guidelines, and execute appropriate API calls to reach a target database state.
We evaluate on the Retail domain, which contains realistic customer service scenarios. Success requires: (1) correctly invoking APIs with write capabilities (e.g., cancel_order, update_shipping) to modify the database state, (2) adhering to domain-specific policies, and (3) managing multi-turn interactions to gather necessary information from users. The benchmark measures end-to-end task completion by comparing the final database state with the ground-truth goal state, directly assessing whether improved tool selection translates to successful task completion in realistic scenarios.
We fix the user simulator to Llama-3.1-8B-Instruct and use temperature 0 for deterministic evaluation, reporting the mean pass rate over 5 independent runs (N=5) to account for stochasticity in user simulation. The detailed prompt given to the agent is shown in Appendix 11.
D.5 PA-Tool Integration
Within these evaluation protocols, PA-Tool operates by replacing original tool names and parameter names in the provided schemas with pretraining-aligned alternatives. Critically, all other components remain identical: tool descriptions, parameter descriptions, example demonstrations, and task instructions are unchanged. This ensures that PA-Tool’s improvements stem solely from schema alignment, not from modifications to semantic information or task structure.
For instance, if the original schema contains get_user_id but PA-Tool determines that get_customer_id better aligns with the model’s pretrained knowledge, only the tool name in the candidate list (MetaTool) or tool schema (RoTBench) is replaced. The tool’s description, available parameters, and all contextual information remain identical across Base and PA-Tool conditions, ensuring fair and comparable evaluation. If get_user_id is the correct answer, methods that do not use schema renaming should predict get_user_id to be considered correct, while PA-Tool should choose get_customer_id.
Figure 14: Average performance gain (diff: PA-Tool - Base) across different models and metrics. (Top) MetaTool results. (Bottom) RoTBench results.
Appendix E Generalization to Diverse Models Details
To assess whether PA-Tool generalizes across different model families and scales, we conduct additional experiments on MetaTool and RoTBench using models beyond the Qwen and Llama families examined in our main results.
For small language models (SLMs), we evaluate three additional models: Ministral-8B, GPT-4.1-nano, and Gemini-2.5-Flash-Lite. To investigate whether our approach remains effective for larger models, we also test on Llama-3.3-70B, GPT-4.1-mini, and Gemini-2.5-Flash. Figure 14 presents the average performance gains across all evaluation metrics for each model. We present detailed performance gains in Figures 12 and 13 for MetaTool and RoTBench, respectively.
Results on SLMs.
Consistent with our main findings, PA-Tool substantially improves performance over the Base setting across all SLMs. Notably, on RoTBench, PA-Tool achieves performance gains exceeding 11% for all three models, demonstrating significant improvements when SLMs serve as executors for subtasks. These results reinforce our key insight: aligning tool call schemas with the model’s intrinsic knowledge provides substantial benefits, even for resource-constrained models.
Results on LLMs.
Our method also yields consistent improvements on larger models. For the closed-source models GPT-4.1-mini and Gemini-2.5-Flash, we observe notable gains of over 9.9% on MetaTool and 10.4% on RoTBench. Interestingly, Llama-3.3-70B shows more modest improvements, suggesting that some larger open-source models may already possess sufficient capability to handle misaligned schemas. Nevertheless, the consistent gains across most models indicate that PA-Tool provides value across different model scales.
These results demonstrate that PA-Tool is a general-purpose technique applicable across diverse model families, architectures, and scales, rather than being specific to particular models or model sizes.
Appendix F Experimental Setup of Supervised Fine-tuning
Setup. As MetaTool does not include a predefined training split, we construct train/validation/test sets by randomly sampling 60%/20%/20% of all task instances. Since the dataset only provides ground-truth tool labels without reasoning traces, we use GPT-4.1-mini to generate complete responses with reasoning for each training instance. Using these generated trajectories, we train our SFT1 baseline. To investigate whether PA-Tool remains effective even when models are trained on larger datasets that naturally reduce format inconsistencies, we create an augmented training set by doubling the original data and train an additional SFT2 baseline. The prompt template is provided in Appendix B. We fine-tune Llama-3.1-8B using the hyperparameters detailed in Table 4.
Table 4: Training hyperparameters for Llama3.1-8B fine-tuning.
Hyperparameter Value
Base Model Llama3.1-8B
LoRA Rank 32
LoRA Alpha 64
LoRA Dropout 0.05
Learning Rate 5e-5
Batch Size 16
Epochs 5
GPU NVIDIA A6000 D6 48GB
Appendix G Error Analysis Details
Error taxonomy. We categorize tool selection errors into three types based on their underlying failure modes: (1) Schema Misalignment Error occurs when the model generates a plausible but nonexistent tool name, or selects a tool with similar naming patterns but different from the correct one. This indicates the model follows its pretrained conventions rather than the provided schema. For example, when the correct tool is get_user_id, the model might generate get_customer_id—a naming pattern it encountered frequently during pretraining but absent from the current schema. (2) Functional Confusion Error occurs when the model selects an existing tool with similar functionality to the correct tool but is not the right choice. This suggests the model understands the query but confuses functionally similar tools. For instance, when asked to send an email notification, the model might select send_sms instead of the correct send_email, both of which are communication tools but use different channels. (3) Context Understanding Error occurs when the model fails to comprehend the query’s intent, selecting a functionally unrelated tool. For example, when asked to delete a user account, the model might incorrectly select create_user or list_products, indicating a fundamental misunderstanding of the task.
Experimental setup. We implement an automated error analyzer using GPT-4.1-mini OpenAI (2024) to classify errors from Llama3.1-8B in both Base and PA-Tool configurations. The analysis is conducted on MetaTool’s three tool selection subtasks: Similar, Scenario, and Multi-tool. We exclude the Reliability subtask as it specifically tests the "no suitable tool" scenario rather than tool selection errors. The classification prompt is provided in Figure 6.
Model MetaTool RoTBench
Qwen2.5-3B 8.0 sec 13.9 sec
Qwen2.5-7B 11.2 sec 15.0 sec
Llama3.2-3B 8.2 sec 12.4 sec
Llama3.1-8B 11.6 sec 15.9 sec
Table 5: Schema generation time for PA-Tool.
Model MetaTool RoTBench
Qwen2.5-3B 2 (1.0%) 13 (4.2%)
Qwen2.5-7B 0 (0.0%) 15 (4.8%)
Llama3.2-3B 1 (0.5%) 18 (5.8%)
Llama3.1-8B 1 (0.5%) 17 (5.5%)
Table 6: Name collision statistics across benchmarks.
Appendix H Computational Time
Table 5 shows the wall-clock time required for one-time schema generation across both benchmarks.
Schema generation is a one-time preprocessing cost, requiring 8–16 seconds depending on model size and number of tools. Once generated, the pretraining-aligned schema can be reused indefinitely without additional overhead during inference. All experiments were conducted on Intel Xeon Gold 6230 CPU @ 2.10GHz (38 cores), NVIDIA A100 80GB PCIe, 450GB RAM, with Python 3.11.14, vLLM 0.10.2, and CUDA 12.8.
Appendix I Name Collision Resolution
When tools have highly similar descriptions, PA-Tool may generate identical names for different components. We resolve this through iterative priority-based locking: in each round, the component with the highest peakedness for a contested name acquires it, while others cascade to their next candidate. This process repeats until all components have unique names.
Table 6 shows collision frequencies across benchmarks. Collisions remain rare across all models (less than 6%), with MetaTool showing particularly low rates (0-1.0%) and RoTBench showing slightly higher but still modest rates (4.2-5.8%). RoTBench’s higher collision rate stems from greater lexical overlap among tool descriptions (average Jaccard similarity of 0.025 vs. 0.011 for MetaTool).
Table 7: Effect of Number of Candidates on Model Performance
Model N MetaTool
Similar Scenario Reliab. Multi.
Qwen2.5-3B 16 50.5 59.1 85.7 66.4
32 50.0 58.8 86.2 72.6
64 50.9 59.1 86.1 69.6
128 50.5 58.6 86.0 70.6
256 50.2 59.1 83.2 72.4
Qwen2.5-7B 16 63.2 74.9 86.8 84.3
32 64.1 78.4 88.2 84.9
64 62.9 77.7 88.2 87.7
128 62.3 77.4 85.4 84.1
256 63.0 78.4 81.0 87.1
Llama3.2-3B 16 64.2 65.9 60.4 80.9
32 65.7 67.7 60.6 80.5
64 64.3 65.2 57.5 79.5
128 63.7 65.2 52.5 79.3
256 64.8 65.5 54.2 80.7
Llama3.1-8B 16 68.9 79.5 66.9 88.9
32 70.3 79.9 66.0 88.3
64 68.7 77.8 64.5 87.5
128 67.9 79.4 65.1 88.9
256 67.6 79.8 63.9 88.3
Appendix J Hyperparameter Detailed Results
Figure 15: Impact of hyperparameters on PA-Tool across different models. All results are averaged across four MetaTool subtasks. Top: Effect of the number of candidates (NN). Middle: Effect of similarity threshold ($\alpha$). Bottom: Effect of sampling temperature ($t$).
We investigate the impact of PA-Tool’s three key hyperparameters on performance: the number of candidates ($N$), the similarity threshold ($\alpha$), and the sampling temperature ($t$). All experiments are conducted on the MetaTool benchmark.
Number of Candidates ($N$).
Table 7 shows that smaller models (3B) achieve stable performance with 16–32 candidates, while larger models (7–8B) require 32–64 candidates before performance plateaus. Beyond these ranges, additional candidates provide minimal gains.
Similarity Threshold ($\alpha$).
Table 8 shows that performance peaks at $\alpha=0.2$ consistently across all models. At $\alpha=0.1$, performance drops by 2–3%, while $\alpha=0.3$ shows similar degradation.
Sampling Temperature ($t$).
Table 9 shows that performance remains stable across temperatures $t\in[0.2,1.0]$, varying within 1–2%. Moderate temperatures (t=0.4t=0.40.60.6) show slightly better results, though the differences are marginal.
Table 8: Effect of Similarity Threshold on Model Performance
Model $\alpha$ MetaTool
Similar Scenario Reliab. Multi.
Qwen2.5-3B 0.1 50.5 58.3 84.1 72.0
0.2 50.0 58.8 86.2 72.6
0.3 51.0 58.6 78.9 60.2
Qwen2.5-7B 0.1 62.7 76.4 84.8 84.9
0.2 64.1 78.4 88.2 84.9
0.3 63.2 75.8 84.7 82.3
Llama3.2-3B 0.1 62.8 66.3 54.8 68.2
0.2 65.7 67.7 60.6 80.5
0.3 65.0 65.5 59.0 79.9
Llama3.1-8B 0.1 68.3 76.9 65.0 88.1
0.2 70.3 79.9 66.0 88.3
0.3 68.1 78.7 64.4 88.7
Appendix K Integration with Training-free Methods Details
K.1 Retrieval-based Correction
Retrieval-based methods attempt to recover from schema misalignment by mapping invalid model outputs to valid tool names post-hoc.
BM25.
When the model generates a tool name not present in the predefined schema, we retrieve the most similar valid tool name using BM25 scoring (Robertson and Zaragoza, 2009) based on character-level lexical overlap. This approach assumes that misaligned outputs share surface-level similarity with intended tools.
ToolLLM Embedding.
We employ ToolLLM’s tool-specialized embedding model (Qin et al., 2024), which was trained specifically for tool retrieval tasks. When misalignment occurs, we compute semantic similarity between the generated output and all valid tool names, selecting the nearest neighbor in embedding space.
Correction Mechanism.
Both methods operate as post-hoc recovery applied after model inference. Specifically, we first evaluate each prediction using the standard scoring procedure. For incorrect predictions, we parse the model output to extract the predicted tool name. If this name does not appear in the predefined candidate list (checked case-insensitively), we apply retrieval-based correction to find the nearest valid tool. The corrected output then undergoes re-evaluation. Crucially, if the predicted tool is present in the candidate list but is simply the wrong choice, no correction is applied—retrieval methods can only recover from schema-violating outputs, not from semantically incorrect selections among valid options.
Results.
As shown in Table 11, retrieval-based correction yields limited improvements across both benchmarks. On MetaTool, BM25 and ToolLLM produce marginal or no gains in most subtasks, with accuracy improvements rarely exceeding 1%. RoTBench shows slightly more responsiveness to retrieval, particularly for tool selection in single-turn settings, but parameter identification remains largely unaffected. Notably, when combined with PA-Tool, retrieval methods achieve more consistent improvements, suggesting that better-aligned initial generations provide more suitable candidates for retrieval-based matching.
Table 9: Effect of Sampling Temperature on Model Performance
Model Temp. MetaTool
Similar Scenario Reliab. Multi.
Qwen2.5-3B 0.2 50.2 58.5 85.0 71.0
0.4 50.0 58.8 86.2 72.6
0.6 51.5 58.9 85.5 70.2
0.8 50.5 59.7 83.9 72.0
1.0 51.7 60.2 85.1 67.2
Qwen2.5-7B 0.2 62.2 75.3 88.8 86.9
0.4 64.1 78.4 88.2 84.9
0.6 62.3 75.8 88.6 86.9
0.8 61.4 77.7 87.5 86.1
1.0 63.1 77.4 90.2 86.7
Llama3.2-3B 0.2 64.1 68.3 55.6 74.2
0.4 65.7 67.7 60.6 80.5
0.6 61.3 66.4 64.6 79.9
0.8 63.5 68.7 54.1 80.3
1.0 64.0 66.3 61.5 77.9
Llama3.1-8B 0.2 69.5 79.5 63.5 88.5
0.4 70.3 79.9 66.0 88.3
0.6 68.4 79.8 62.3 80.9
0.8 67.8 79.4 65.2 88.7
1.0 69.5 76.9 65.6 88.5
Model Method MetaTool RoTBench
Tool Selection Single-turn Multi-turn
Similar Scenario Reliability Multi-tool Tool Sel. Param Iden. Tool Sel. Param Iden.
GPT-4.1-mini Base 79.6 84.3 76.3 72.2 79.1 58.1 28.6 12.9
PA-Tool 80.5 84.5 80.0 84.3 82.9 59.1 25.7 21.4
GPT-4.1-nano Base 56.8 53.1 99.0 81.5 71.4 34.3 28.6 12.9
PA-Tool 60.6 54.0 99.5 77.1 72.4 44.8 27.1 17.1
Gemini-2.5-Flash Base 70.0 79.8 89.2 77.3 82.9 56.2 22.9 12.9
PA-Tool 72.6 82.7 92.4 78.5 84.8 62.9 31.4 17.1
Gemini-2.5-Flash-Lite Base 79.3 86.5 65.9 77.7 72.4 63.8 24.3 18.6
PA-Tool 80.8 87.1 70.5 72.0 73.3 65.7 30.0 22.9
Llama3.3-70B Base 80.4 85.2 63.7 82.9 75.2 27.6 32.9 8.6
PA-Tool 79.9 85.2 66.2 81.3 74.3 26.7 30.0 7.1
Ministral-8B Base 65.3 81.2 42.5 93.1 66.7 28.6 31.4 15.7
PA-Tool 70.5 82.6 44.7 92.4 75.2 31.4 31.4 15.7
Table 10: Performance comparison on MetaTool and RoTBench with additional models.
Model Method MetaTool RoTBench
Tool Selection Single-turn Multi-turn
Similar Scenario Reliability Multi-tool Tool Sel. Param Iden. Tool Sel. Param Iden.
Qwen2.5-3B Base 48.7 55.3 83.6 75.1 1.9 0.0 1.4 0.0
BM25 48.7 55.3 83.6 75.1 3.8 0.9 4.3 0.0
ToolLLM 48.7 55.3 83.7 75.1 4.8 1.9 7.1 2.9
PA-Tool 49.9 58.8 86.2 72.6 6.7 1.9 7.1 4.3
PA-Tool + BM25 50.3 60.2 86.2 72.8 10.5 2.9 10.0 7.1
PA-Tool + ToolLLM 50.3 60.3 86.2 73.4 10.5 3.8 12.9 10.0
Qwen2.5-7B Base 59.6 74.4 78.3 78.3 49.5 20.0 8.6 5.7
BM25 59.8 74.6 78.3 78.3 49.5 20.9 8.6 5.7
ToolLLM 59.8 74.6 78.3 78.3 49.5 20.9 12.9 8.6
PA-Tool 64.1 78.4 88.2 84.9 55.2 21.9 18.6 11.4
PA-Tool + BM25 64.1 78.5 88.2 84.9 56.2 21.9 18.6 11.4
PA-Tool + ToolLLM 64.1 78.5 88.2 85.5 55.2 21.9 18.6 11.4
Llama3.2-3B Base 55.0 58.6 43.6 79.1 56.2 20.0 27.1 5.7
BM25 55.7 58.7 43.6 79.1 56.2 22.9 27.1 4.3
ToolLLM 55.4 58.8 43.6 79.1 56.2 22.9 27.1 4.3
PA-Tool 65.7 67.7 60.6 80.5 62.9 21.9 31.4 5.7
PA-Tool + BM25 65.7 67.7 60.6 80.5 63.8 22.9 31.4 5.7
PA-Tool + ToolLLM 65.7 67.7 60.6 80.5 62.9 21.9 31.4 4.3
Llama3.1-8B Base 61.5 73.9 53.5 78.7 58.1 17.1 17.1 4.3
BM25 61.8 74.0 53.5 78.7 58.1 19.1 17.1 2.9
ToolLLM 61.6 74.0 53.5 78.7 58.1 19.1 17.1 2.9
PA-Tool 70.4 79.9 66.0 88.3 68.6 18.1 22.9 4.3
PA-Tool + BM25 70.4 79.9 66.0 88.3 68.6 18.1 22.9 2.9
PA-Tool + ToolLLM 70.4 79.9 66.0 88.3 68.6 19.1 22.9 2.9
Table 11: Performance of retrieval-based post-hoc correction methods (BM25, ToolLLM) on MetaTool and RoTBench.
Model Method MetaTool RoTBench
Tool Selection Single-turn Multi-turn
Similar Scenario Reliability Multi-tool Tool Sel. Param Iden. Tool Sel. Param Iden.
Qwen2.5-3B Base 48.7 55.3 83.6 75.1 1.9 0.0 1.4 0.0
PA-Tool 50 58.8 86.2 72.6 6.7 1.9 7.1 4.3
Constrained 65.9 80.3 60.9 77.3 65.7 16.2 20.0 4.3
Constrained-PA 66.1 77.9 74.5 76.5 61.9 26.7 25.7 8.6
Qwen2.5-7B Base 59.6 74.4 78.3 78.3 49.5 20 8.6 5.7
PA-Tool 64.1 78.4 88.2 84.9 55.2 21.9 35.7 11.4
Constrained 64.2 81.0 78.4 83.3 78.1 31.4 32.9 10.0
Constrained-PA 67.9 79.3 84.0 84.1 76.2 31.4 34.3 12.9
Llama3.2-3B Base 55.0 58.6 43.6 79.1 56.2 20.0 27.1 5.7
PA-Tool 65.7 67.7 60.6 80.5 62.9 21.9 31.4 5.7
Constrained 55.1 70.2 72.3 89.5 62.9 20.0 25.7 5.7
Constrained+PA-Tool 59.2 69.8 65.5 86.9 68.6 21.9 28.6 4.3
Llama3.1-8B Base 61.5 73.9 53.5 78.7 58.1 17.1 17.1 4.3
PA-Tool 70.4 79.9 66.0 88.3 68.6 18.1 22.9 4.3
Constrained 66.3 81.7 66.2 74.3 64.8 16.2 25.7 5.7
Constrained+PA-Tool 69.8 81.2 66.3 79.9 63.8 20.0 30.0 4.3
Table 12: Performance of Constrained generation using JSON Schema enum on MetaTool and RoTBench
K.2 Constrained Generation
Constrained generation methods enforce valid outputs during the decoding process itself, eliminating the possibility of schema-violating generations.
JSON Schema Constraints.
We implement constrained decoding using JSON Schema with enum restrictions, which masks logits corresponding to tokens that would lead to invalid tool names during generation. For MetaTool, the schema restricts outputs to {"selected_tool": <enum>}, while for RoTBench it enforces {"action": <enum>, "action_input": <object>}, where the enum contains all valid tool names plus special tokens (e.g., None, finish).
Results.
As shown in Table 12, constrained generation using JSON Schema enum substantially improves performance on RoTBench by eliminating format errors, with Qwen2.5-7B achieving 78.1% on single-turn tool selection compared to 49.5% with the base model. However, results on MetaTool are more mixed, with constrained generation sometimes underperforming PA-Tool alone, particularly on the Reliability subtask. Combining constrained generation with PA-Tool (Constrained-PA) yields complementary benefits: the approach maintains the format guarantees of constrained decoding while leveraging PA-Tool’s improved schema alignment. This combination achieves the best results on parameter identification tasks across most models, suggesting that PA-Tool’s aligned representations help the model make better selections even when the output space is explicitly constrained.
Model Method MetaTool RoTBench
Tool Selection Single-turn Multi-turn
Similar Scenario Reliability Multi-tool Tool Sel. Param Iden. Tool Sel. Param Iden.
Qwen2.5-3B Base 48.7 55.3 83.6 75.1 1.9 0.0 1.4 0.0
PA-Tool 50 58.8 86.2 72.6 6.7 1.9 7.1 4.3
EasyTool 52.3 64.3 91.5 70.6 13.3 9.5 4.3 1.4
EasyTool+PA-Tool 48.0 47.7 93.4 46.1 20.0 11.4 10.0 5.7
Qwen2.5-7B Base 59.6 74.4 78.3 78.3 49.5 20 8.6 5.7
PA-Tool 64.1 78.4 88.2 84.9 55.2 21.9 35.7 11.4
EasyTool 65.1 78.9 83.0 85.1 50.5 26.7 12.9 7.1
EasyTool+PA-Tool 64.5 73.5 90.5 79.3 52.4 26.7 17.1 10.0
Llama3.2-3B Base 55.0 58.6 43.6 79.1 56.2 20.0 27.1 5.7
PA-Tool 65.7 67.7 60.6 80.5 62.9 21.9 31.4 5.7
EasyTool 45.3 65.3 64.1 79.5 56.2 20.9 27.1 5.7
EasyTool+PA-Tool 42.0 58.3 80.5 66.6 66.7 23.8 28.6 4.3
Llama3.1-8B Base 61.5 73.9 53.5 78.7 58.1 17.1 17.1 4.3
PA-Tool 70.4 79.9 66.0 88.3 68.6 18.1 22.9 4.3
EasyTool 61.1 77.2 76.9 72.2 59.0 14.3 21.4 2.9
EasyTool+PA-Tool 66.2 77.9 73.8 68.6 64.8 21.9 20.0 2.9
Table 13: Performance of description enhancement (EasyTool) on MetaTool and RoTBench
K.3 Description Enhancement
Description enhancement methods improve tool understanding by rewriting documentation for clarity, without modifying the tool naming schema.
EasyTool.
EasyTool (Yuan et al., 2025) creates unified tool instructions by restructuring verbose API documentation into clear, consistent formats. The original method uses GPT-4 to transform tool descriptions into structured representations containing concise functionality summaries, enumerated API lists, and usage examples.
We adapt this methodology with two key modifications. First, to ensure fair comparison without introducing external model capabilities, we use the target model itself to generate improved descriptions rather than GPT-4. Second, we simplify the output format to focus on description clarity: rather than creating full structured instructions, we generate concise 1-2 sentence descriptions using action-oriented verbs (e.g., “Retrieves”, “Calculates”) while explicitly removing tool name references to prevent self-referential artifacts. The prompt template follows EasyTool’s core principles of removing redundancy and improving clarity (Figure 16).
Results.
As shown in Table 13, PA-Tool consistently outperforms EasyTool across most configurations. The performance gap is particularly notable on MetaTool, where Llama3.2-3B achieves 65.7% on the Similar subtask with PA-Tool compared to only 45.3% with EasyTool. Similarly, Llama3.1-8B shows substantial improvements with PA-Tool over EasyTool on both Similar (70.4% vs. 61.1%) and Multi-tool (88.3% vs. 72.2%) subtasks. These results suggest that naming alignment is more effective than description enhancement for improving tool selection. Furthermore, rewriting descriptions carries the risk of altering the perceived functionality of tools, potentially introducing semantic drift. In contrast, PA-Tool preserves the original semantics by modifying only the tool names. Nevertheless, the two approaches remain compatible, with some combined configurations (EasyTool+PA-Tool) showing competitive results, particularly on the Reliability subtask where the combination achieves the highest scores across all models.
Description Enhancement Prompt (EasyTool-style) You are tasked with improving a tool description following the EasyTool methodology. The goal is to create clear, structured, and unified instructions that help language models better understand the tool’s purpose and usage. Original Description: {{ description }} Instructions: 1. Create a CONCISE description that clearly states the tool’s core functionality 2. Focus on WHAT the tool does and WHEN to use it 3. Remove redundant, verbose, or marketing-style language 4. Use action-oriented verbs at the start (e.g., “Retrieves”, “Calculates”, “Searches”) 5. Keep the description to 1-2 sentences maximum Critical Rules: Do NOT include ANY tool name, API name, or brand name Do NOT start with “This tool…”, “The tool…”, or any name reference Start DIRECTLY with an action verb describing the functionality Example: Original: “WeatherAPI is a comprehensive weather tool that can be used to get current weather data and forecasts.” Improved: “Retrieves current weather conditions and forecasts including temperature, humidity, and precipitation for specified locations.” Generate ONLY the improved description without additional explanation.
Figure 16: Prompt template for description enhancement following EasyTool (Yuan et al., 2025) methodology. Unlike the original work which uses GPT-4 to create structured unified instructions, we use the target model itself to generate concise descriptions, ensuring fair comparison.